### v1

# Who Gets Cited Most? Benchmarking Long-Context Language Models on Scientific Articles

Abstract
This paper introduces SciTrek, a novel question-answering benchmark designed to evaluate the long-context reasoning capabilities of large language models (LLMs) using scientific articles. Current long-context benchmarks often rely on non-scientific texts, focus on simple information retrieval tasks, or employ artificial contexts. SciTrek addresses these limitations by proposing complex questions that require information aggregation and synthesis across multiple full-text scientific articles. Questions and their ground-truth answers are automatically generated by formulating them as SQL queries over a database constructed from article metadata (titles, authors, and references). The SQL operations provide explicit, verifiable reasoning steps for fine-grained error analysis, and the construction process scales to contexts up to 1M tokens with minimal supervision. Extensive experiments on a diverse set of open-weight and proprietary LLMs demonstrate that SciTrek poses a significant challenge as the context length increases, with supervised fine-tuning and reinforcement learning offering only limited gains. Our analysis reveals systematic shortcomings in models’ abilities to perform basic numerical operations and accurately locate specific information in long contexts.111Our dataset can be downloaded from https://github.com/oaimli/SciTrek
1 Introduction
Large language models (LLMs) show promise in accelerating scientific progress by assisting researchers in reviewing and synthesizing the growing body of literature (Pearson, 2024; Agarwal et al., 2025) and generating novel research ideas (Si et al., 2025; Baek et al., 2025). So much so that industry products such as Deep Research222https://openai.com/index/introducing-deep-research/, Elicit333https://elicit.com/, and Scite444https://scite.ai/ have emerged as dedicated tools to aid users with complex, multi-step research tasks. Scientific workflows often require processing large inputs: full scientific articles, collections of references, or structured datasets. A model capable of processing 100K+ tokens could simultaneously analyze multiple articles, track long chains of reasoning, or connect experimental results to prior work.
While long-context language models (LCLMs) are highly relevant to scientific tasks, existing benchmarks designed to evaluate their capabilities predominantly focus on non-scientific texts (Maekawa et al., 2025; Bai et al., 2025; Yen et al., 2025; Kuratov et al., 2024). They mostly address simple information retrieval tasks (e.g., Needle-In-A-Haystack; Kamradt 2023; Hsieh et al. 2024) rather than information integration, i.e., reasoning across multiple documents to synthesize information and answer complex questions. Some question-answering benchmarks (Asai et al., 2024; Bai et al., 2025; An et al., 2024) require information integration, but do not provide a typology of the reasoning steps involved, which makes it difficult to determine the specific capabilities being assessed and why a model fails. Another issue relates to scalability and extensibility. When substantial human effort is involved in generating the questions and their answers (e.g., Asai et al. 2024), scaling to longer contexts or different types of questions becomes prohibitive. To better emulate scientific research practice, it is also important to evaluate a range of skills in natural contexts, such as a model’s ability to aggregate or filter information or combinations thereof (Wang et al., 2024c). However, a few benchmarks rely on artificial contexts, often created by extending short-context tasks (e.g., HotpotQA; Yang et al. 2018) with added noise (Kuratov et al., 2024) or by synthesizing the input with LLMs (Maekawa et al., 2025).
In this paper, we create a question-answering benchmark based on scientific articles that aims to alleviate these issues. We assume a scenario where a researcher, possibly during the literature review process, seeks answers to specific questions about the scientific articles they read. As an example, consider the first question in Table 1: “What is the highest number of authors that any single article has?”. Although seemingly simple, it requires aggregation of information across multiple articles. To answer correctly, a hypothetical model would need to count the number of authors for each article within a document collection, and identify the article with the largest number. To generate questions with natural language inputs and explicit reasoning steps, we focus on basic numerical operations such as counting, aggregating, and sorting, applied to key elements of scientific articles: titles, authors, and references. Since these operations correspond to standard database functions, we can automatically generate arbitrarily complex SQL queries over any collection of scientific articles and obtain ground-truth answers without manual annotation by constructing database tables for titles, authors, and references. We take advantage of recent advances in LLM capabilities for SQL understanding and generation (Yu et al., 2018; Li et al., 2023; Hui et al., 2024) to automatically convert SQL queries into corresponding natural language questions.
| Question | SQL Query | Answer |
| What is the highest number of authors that any single article has? | SELECT MAX(author_count) FROM articles |
10
| What is the word count of the titles of articles, sorted by the number of authors in ascending order? | SELECT title_word_count FROM articles ORDER BY author_count ASC | 9, 17, 5, 6, 9, 12 |
| How many references do articles with exactly two authors have? | SELECT reference_count FROM articles WHERE author_count = 2 | 16 |
| What is the total number of words in the titles of all articles that have exactly 60 references? | SELECT SUM(title_word_count) FROM articles WHERE reference_count 60 |
13
| What are the names of the authors who are either the first or second author of an article, listed in descending order of their position? | SELECT author_name FROM article_author WHERE author_position < 2 ORDER BY author_position DESC | N. Shazeer, A. Vaswani |
| How many articles have been cited by other articles but do not cite any other articles? | SELECT COUNT(*) FROM articles WHERE article_id NOT IN (SELECT article_id_citing FROM citing_cited) AND article_id IN (SELECT article_id_cited FROM citing_cited) |
2
Our benchmark, SciTrek, contains a test set of 2,121 question-answer pairs on scientific articles and a training set of 19,543 question-answer pairs, with contexts of varying lengths (i.e., 64K, 128K, 512K, and 1M555Throughout the paper, K refers to 1,024 and M to 1,024K.). Table 1 illustrates examples of questions, their corresponding SQL queries, and answers. These questions are relatively superficial, as they do not engage deeply with article content. However, the metadata consists of distillable facts sparsely distributed across article collections, requiring models to retrieve and synthesize information over long contexts. If models struggle with simple filtering and sorting, they are unlikely to succeed on more complex analytical tasks. We use SciTrek to evaluate the capabilities of LCLMs with context windows exceeding 128K tokens. Experimental results reveal that the benchmark poses significant challenges, with models achieving poor performance, even when enhanced with supervised fine-tuning (SFT) and reinforcement learning (RL). We make the following contributions:
We propose a methodology for constructing a long-context question-answering benchmark over natural texts, with explicit reasoning steps approximated via SQL operations that involve information aggregation. The methodology requires minimal human intervention and can be replicated for other domains containing entities that can be structured as database tables.
Extensive experiments reveal that frontier open-weight and proprietary LLMs struggle significantly with this task, especially as input length increases. While SFT and RL improve over baseline models, performance remains limited. All models struggle with questions involving references and citation relationships among the input articles.
By design, SciTrek enables fine-grained analysis of model behavior, offering insight into where and why models fail. Our evaluation of open-weight models reveals systematic shortcomings in counting bibliographic elements and performing basic numerical operations, with error rates increasing as question length grows. Interestingly, models frequently misinterpret compound conditions and struggle with logical constructs involving negation. (see the last question in Table 1).
| Dataset | Natural | Skills | Reason | Scalable | Length | Example Question |
| NeedleBench | ✗ | ✗ | ✗ | ✓ | 128K | What legendary item is hidden on Emerald Island? |
| Ada-LEval | ✗ | ✗ | ✗ | ✓ | 128K | What is the correct order of the segments? |
| BABILong | ✗ | ✗ | ✗ | ✓ | 10M | Where is Mary? |
| HELMET | ✗ | ✓ | ✗ | ✗ | 128K | Who set the fire in one tree hill? |
| LIFBENCH | ✗ | ✓ | ✗ | ✓ | 128K | Retrieve the entry at position 8th in the list. |
| RULER | ✗ | ✓ | ✗ | ✓ | 128K | Find all variables that are assigned the value 12345. |
| OpenScholar | ✓ | ✓ | ✗ | ✗ | 3K∗ | Compile a list of reviews […], and identify the most promising […] |
| LongBench v2 | ✓ | ✓ | ✗ | ✗ | 128K | How long have I been living in my current apartment in Shinjuku? |
| LongMemEval | ✓ | ✓ | ✗ | ✗ | 2M | How many bikes do I currently own? |
| L-Eval | ✓ | ✓ | ✗ | ✗ | 200K | How do I know when I should apply for Medicare? |
| HoloBench | ✗ | ✓ | ✓ | ✓ | 64K | What are the names of wines and their corresponding grape types? |
| MathHay | ✗ | ✓ | ✓ | ✓ | 128K | What is the total number of points scored by LeBron […] combined? |
| Loong | ✓ | ✗ | ✓ | ✓ | 250K | Which company has the highest non-current assets? |
| SciTrek | ✓ | ✓ | ✓ | ✓ | 1M | How many articles have been cited by other articles but do not cite any other articles? |
2 Related Work
Long-context Language Modeling
A growing body of work explores the challenges LLMs encounter when processing inputs that exceed their context lengths. Rotary Position Embeddings (RoPE; Su et al. 2024) have emerged as a widely used approach, replacing absolute positional embeddings with rotational transformations of token embeddings. It improves generalization to unseen sequence lengths. Building on RoPE, several extensions introduce rescaling to further enhance performance on substantially longer inputs without the need for retraining (Peng et al., 2024).
Other architectural advances focus on reducing the quadratic cost of the Transformer attention mechanism (Sun et al., 2025). Sparse attention methods, such as LongFormer (Beltagy et al., 2020; Jiang et al., 2024; Fu et al., 2024), compute only selected portions of the full attention matrix, thereby reducing computational cost and accelerating both prefilling and inference. Many contemporary models interleave sparse and full attention layers to balance efficiency and performance (Llama Team, 2025; Yang et al., 2025; Kamath et al., 2025; Dubey et al., 2024). Another popular strategy of reducing the computation required in long-context settings is to limit the number of active model parameters. Group Query Attention (GQA; Ainslie et al. 2023) enables multiple attention heads to share key–value projections, without significant performance loss. The open-weight LLMs used in our experiments adopt combinations of these techniques, complemented by pretraining on long-context tasks (Dubey et al., 2024; Kamath et al., 2025; Yang et al., 2025; Llama Team, 2025).
Long-context Datasets
With growing interest in LCLMs, there has been a surge in benchmarks designed to evaluate their performance. As far as text-based efforts are concerned, initial benchmarks primarily assessed whether models are able to retrieve relevant information from their context, typically through Needle-In-A-Haystack (NIAH) tasks, such as NeedleBench (Li et al., 2024). Building on these, subsequent benchmarks introduced either distractor information to short-context tasks or dispersed relevant information across synthetic contexts, such as Ada-LEval (Wang et al., 2024a), BABILong (Kuratov et al., 2024), HELMET (Yen et al., 2025), LIFBENCH (Wu et al., 2025b), LongBench (Bai et al., 2024), and RULER (Hsieh et al., 2024).
In addition to retrieval, some benchmarks emphasize complex reasoning over long contexts. For example, OpenScholar (Asai et al., 2024) and LongBench v2 (Bai et al., 2025) contains questions that require extensive use of the multiple documents provided as context, while DocFinQA (Reddy et al., 2024) and MedOdyssey (Fan et al., 2025) focus on financial and medical reasoning, respectively. Although not explicitly technical, LongMemEval (Wu et al., 2025a) and L-Eval (An et al., 2024) also requires reasoning over natural, long-context information. However, evaluating model answers against these benchmarks remains challenging, or the reasoning steps required to arrive at correct responses are often opaque, making it difficult to assess failure modes of LCLMs. In contrast, HoloBench (Maekawa et al., 2025) and MathHay (Wang et al., 2024b) better delineate between the types of reasoning required for each question, but rely on unnatural contexts. Loong (Wang et al., 2024c) combines structured reasoning and natural contexts, but tests a very limited number of skills. CURIE (Cui et al., 2025), OpenScholar (Asai et al., 2024) and LongBench v2 (Bai et al., 2025) evaluate reasoning over scientific articles, but their reliance on expert-curated questions and answers poses challenges for scalability to longer contexts or larger datasets.
In designing SciTrek, we aim to retain many of the useful qualities of previous benchmarks (e.g., structured reasoning, scalability) while still proposing a fairly natural and easy task for humans to perform. Although we focus on scientific articles, our methodology extends to other domains with explicit entities and relations. Thanks to the SQL backbone, we are able to construct questions testing various model skills (e.g., aggregation, filtering), and how these manifest themselves through explicit reasoning steps. We compare SciTrek with representative existing benchmarks in Table 2.
3 The SciTrek Benchmark
In this section, we explain how SciTrek was curated. As mentioned earlier, our benchmark consists of question-answer pairs over scientific articles. Figure 1 illustrates our process: we first gather scientific article collections as contexts corresponding to different lengths (i.e., 64K, 128K, 512K, and 1M); we then create databases representing the article metadata and SQL queries with answers from database execution; finally, we convert the SQL queries into natural language questions.
3.1 Gathering Scientific Article Collections
SciTrek is constructed using scientific articles from Semantic Scholar666https://www.semanticscholar.org/. To cover diverse topics of articles, we obtain an initial set of seed articles from eight subjects: Computer Science (CS), Economics, Electronic Engineering (EE), Math, Physics, Biology, Finance, and Statistics.777Following the typology of subjects from https://arxiv.org/. For each subject, we select two seed articles with more than 100 citations since 2020. For each article, we retrieve related articles from Semantic Scholar.888Semantic Scholar provides APIs to retrieve scientific articles with metadata including titles, authors, a reference list with articles that the current article cites, and a citation list with articles that cite the current article. Throughout this paper, a reference is a bibliographic entry listed in the reference section of a scientific article, whereas a citation is an in-text acknowledgement of another’s work within the article. To ensure broad coverage, we include two-hop related articles identified via Semantic Scholar’s citation graph based on the reference and citation list. For each seed article, we randomly sample ten first-hop related articles, and for each of these, we further sample five second-hop related articles to form an article cluster. Since full texts are required, we filter out articles without PDFs. This process yields 16 article clusters comprising 662 scientific articles with PDFs across eight subjects (see step I in Figure 1). Finally, we convert the collected PDFs into markdown texts using Marker.999https://github.com/datalab-to/marker
From these clusters, we construct article collections of varying context lengths for our question-answering task. By concatenating the markdown texts, we generate collections with lengths of 64K, 128K, 512K, and 1M tokens.101010We can easily construct collections with arbitrary lengths exceeding 1M tokens. Each collection is initialized with a randomly selected article, and additional articles are added until it reaches a specified length level. We expand the collections using two strategies: (1) random sampling from the clusters, and (2) traversing the citation graph within each cluster (using both depth-first and breadth-first search) to construct collections that preserve citation relations among articles. Each collection contains at least four articles, and no two collections share more than half of their articles. From the 662 scientific articles across 16 clusters, we construct 2,612 article collections spanning all four length levels, including 2,027 generated through random sampling and 585 through graph traversal.
| SQL Commands | |
| Aggregating | MAX, MIN, SUM, AVG, COUNT, DISTINCT |
| Sorting | ORDER BY, ASC, DESC, GROUP BY |
| Filtering | WHERE |
| SQL Operators | |
| Comparison | , , , , , , LIKE |
| Arithmetic | , , , /, % |
| Logical | AND, NOT, OR, BETWEEN, IN |
| Skill | Count | Example Query Template |
| Aggregating | 20 | SELECT MAX(author_count) FROM articles |
| Sorting | 27 | SELECT title_word_count FROM articles ORDER BY author_count ASC |
| Filtering | 107 | SELECT author_name FROM article_author WHERE author_position {author-position} |
| Filtering+ Aggregating | 107 | SELECT SUM(title_word_count) FROM articles WHERE reference_count {reference-count} |
| Filtering+ Sorting | 106 | SELECT author_count FROM articles WHERE title_word_count % 2 1 ORDER BY title_word_count DESC |
| Relational Filtering | 20 | SELECT COUNT(*) FROM articles WHERE article_id NOT IN (SELECT article_id_citing FROM citing_cited) AND article_id IN (SELECT article_id_cited FROM citing_cited) |
3.2 Creating Databases and SQL Queries
Once the article collections are assembled, we construct a database for each (see step II in Figure 1). We focus on a core subset of elements shared across all scientific articles: titles, authors, and references. Based on these key elements, each database comprises three tables: articles, article-author, and citing-cited. The articles table contains metadata such as the title, reference count, and title word count. The article-author table captures author information, including names and their positions in the corresponding author list. The citing-cited table records citation relations among articles. A detailed description of these tables is provided in Appendix A. This information is obtained from Semantic Scholar or derived via simple preprocessing, such as splitting titles and counting words.
The core SQL111111Structured Query Language (SQL) is a standardized language for managing and querying relational databases. commands summarized in Table 4 form the foundation for building SQL queries, often combined with operators (comparison, arithmetic, logical) to create more complex queries (since WHERE is used to filter data, it always works in conjunction with SQL operators). Aside from the basic SQL commands in Table 4 (Aggregating, Sorting, and Filtering), we define composite commands based on their combinations (i.e., Filtering+Aggregating and Filtering+Sorting). Using these, we manually create SQL query templates targeting different topics related to key elements of scientific articles, including Author Count, Author List, Reference Count, Title List, and Title Word Count. This process is illustrated in Figure 1 (right panel) and example templates are shown in Table 4; Note that some templates include placeholders to be instantiated with specific values. Finally, to capture authorship and citation relations, we introduce Relational Filtering and the topics of Author Relation and Citation Relation, which specifically target authorship and citation relations.
We collectively refer to the SQL commands in Table 4 as information processing skills, since they test different information processing capabilities. We have various templates per skill, designed to be applicable across all article collections. For each collection, we randomly select 10 templates and instantiate them with collection-specific values for all placeholders. We then execute these queries against the corresponding database to generate ground-truth answers.
3.3 Converting SQL Queries to Natural Language
We use Qwen2.5-Coder-32B-Instruct (Hui et al., 2024) to convert SQL queries into natural language questions. To ensure queries and questions are meaning preserving, we validate each generated question by converting it back to SQL and verifying that both queries produce identical results when executed against the collection database. For each SQL-collection pair, we repeat this process up to 10 times; if no valid question is obtained, we discard the query for that collection. Using the prompts in Appendix B, we successfully generate natural language questions for 82.9% of SQL-collection pairs. Following this, we obtain 2,121 test questions for evaluating models against the four context lengths defined in Section 3.1, and 19,543 instances for training (see Section 4.2). Table 6 presents descriptive statistics for the SciTrek test set. For each context length, our test partition covers all information processing skills and question topics described in Section 3.2. The distribution of SQL commands and operators is provided in Appendix C.
3.4 Data Quality Validation
| Length | Instances | Articles/C | Words/Q | Words/A |
| 64K | 112 | 4.2 | 14.6 | 6.2 |
| 128K | 728 | 5.7 | 16.9 | 7.9 |
| 512K | 667 | 22.7 | 18.1 | 30.9 |
| 1M | 614 | 46.3 | 18.6 | 63.6 |
| Skill | Agree (%) | Align (%) |
| Aggregating | 85.7 | 85.7 |
| Sorting | 85.0 | 80.0 |
| Filtering | 85.7 | 85.7 |
| Filtering+Aggregating | 95.0 | 85.0 |
| Filtering+Sorting | 89.5 | 89.5 |
| Relational Filtering | 89.5 | 73.7 |
| All | 88.3 | 83.3 |
To validate the quality of SciTrek, we used Prolific121212https://www.prolific.com/ to crow-source annotations by asking human annotators to provide answers to our natural language questions. We randomly sampled 120 instances representing the six skills listed in Table 4 (20 instances per skill). Each instance was independently annotated by three annotators. Annotators were asked to answer the questions using database tables rather than full-text articles, as documents spanning 1M tokens are impractical for humans to review. The metadata in the database tables were sourced from Semantic Scholar or generated through simple preprocessing as explained in Section 3.1, with manual corrections applied when necessary to make them aligned with the collected full-text articles in Markdown. More annotation details are in Appendix D. We measured inter-annotator agreement and the alignment of answers obtained from executing SQL queries on the database with human responses, using exact match. Annotators were considered in agreement if two or more provided the same answer to a question. Similarly, our answer was considered aligned with the humans if it matched the answer of two or more annotators. Table 6 reports the average agreement and alignment across 120 instances, showing that human annotators largely agree with each other and that our database-based answers are highly consistent with human responses.
| Full-text Articles | Database TablesΔ | ||||||||
| Model | CLen | 64K | 128K | 512K | 1M | 64K | 128K | 512K | 1M |
| Qwen2.5-7B-Instruct-1M | 1M | 4.5 | 2.8 | 0.3 | 0.0 | 20.5 | 14.3 | 5.0 | 2.0 |
| Qwen2.5-14B-Instruct-1M | 1M | 8.3 | 6.5 | 1.6 | 0.1 | 33.3 | 27.2 | 11.0 | 5.9 |
| Gemma-3-27B-IT | 128K | 6.2 | 3.4 | — | — | 31.8 | 25.0 | 11.3 | 6.1 |
| Llama-4-Scout-17Bx16E-Instruct | 10M | 5.4 | 2.8 | 1.3 | 1.1 | 28.6 | 19.0 | 7.4 | 4.0 |
| Llama-3.3-70B-Instruct | 128K | 8.3 | 3.5 | — | — | 47.0 | 36.2 | 14.6 | 8.1 |
| DeepSeek-R1-Distill-Llama-70B | 128K | 22.0 | 6.0 | — | — | 83.3 | 74.2 | 56.8 | 42.2 |
| Gemini 2.5 Pro | 1M | 41.7 | 26.0 | 91.7 | 83.5 | 55.4 | 31.5 | ||
| GPT-4.1 | 1M | 21.1 | 11.7 | 3.9 | 2.5 | 69.3 | 53.8 | 24.8 | 16.1 |
| o4-mini | 195K | 61.0 | 46.5 | — | — | 95.2 | 87.8 | 79.4 | 72.6 |
4 Results
In our experiments, we evaluate both proprietary and open-weight models from various families that support contexts exceeding 128K and have shown strong performance on language understanding and mathematical reasoning benchmarks, e.g., MMLU (Hendrycks et al., 2021a) and MATH (Hendrycks et al., 2021b). For each family, we select the largest model we can feasibly run with our resources.131313All our experiments were conducted on 4 NVIDIA HGX H200s. For example, we use Llama-4-Scout-17Bx16E-Instruct instead of Llama-4-Maverick-17B-128E-Instruct, and the distilled variant of DeepSeek-R1. Table 7 provides an overview of the models we consider, most of which were released within the past six months. For detailed model descriptions and settings, refer to Appendix F. As shown, the models vary in parameter scale and supported context length. We evaluate models in two context settings: (1) using the full-text scientific articles within a given collection as context and (2) using only the corresponding database tables. All models in our study are instruction-tuned, and we employ a uniform set of prompts across them, as detailed in Appendix E. Models generate three answers for each question. We assess performance using average exact match and F1, as the expected outputs are factual items with minimal variation, such as specific numbers, author names, or article titles.
4.1 Zero-shot Prompting
Our zero-shot results are summarized in Table 7. Overall, we observe that SciTrek is challenging especially when using the full-text articles as contexts. In this setting, performance drops for all models as the input gets longer. This trend also manifests itself when using database tables as context. Perhaps unsurprisingly, proprietary models significantly outperform open-weight ones. We observe similar tendencies when using F1 as the evaluation metric (see Appendix G).
Figure 2 presents a fine-grained analysis across question topics (e.g., author count, reference count, citation relations), skills (e.g., aggregating, sorting), and subjects (e.g., Economics, Biology) when using full-text articles as context. As can be seen, model performance shows little variation by subject. However, most models struggle with sorting tasks, while performing better on aggregation. Performance is lowest on citation-related questions (i.e., Citation Relation, Reference Count), and somewhat higher on author- and title-related questions.
4.2 Long-context Post-training
Leveraging the data generation methodology described in Section 1, we curate a substantial training dataset (19,543 instances) across four context lengths (64K, 128K, 512K, and 1M). We use this data to assess whether more training could improve the performance of open-weight models on full-text articles. We experiment with two well-established techniques: supervised fine-tuning (SFT) and reinforcement learning (RL). Because of the high computational cost, for this suite of experiments, we only report results with Qwen2.5-7B-Instruct-1M on data with maximum context of 128K tokens (7,703 training instances). Moreover, we examine whether the post-trained model can generalize along the dimensions of input length, question topics, and information processing skills. We compare in-distribution performance where the model is trained and tested in similar conditions (e.g., on lengths of 64K and 128K) to out-of distribution where the model is tested on an unseen dimension.
Specifically, we train the model on data corresponding to 64K and 128K context lengths and evaluate it both in-distribution (64K, 128K) and out-of-distribution (512K, 1M). For question topic, training uses author- and title-related questions at 64K and 128K, with evaluation covering the same topics in-distribution and citation-related questions out-of-distribution. For skills, training is performed on non-relational questions (Aggregating, Sorting, Filtering, Filtering+Sorting, Filtering+Aggregating), with evaluation including these in-distribution skills and the out-of-distribution Relational Filtering skill. For SFT, we train the model for 500 steps with a batch size of 32, a learning rate of , and a warm-up rate of 0.05. For RL, we use GRPO (Shao et al., 2024) with a mixed reward of EM and F1 to encourage the model to produce both reasoning traces and answers. GRPO was chosen due to its success in similar verifiable and long-context tasks (Shao et al., 2024; Mroueh, 2025; Gurung & Lapata, 2025; Zheng et al., 2025). We found a simple
| Length | Topic | Skills | ||||
| Models | ID | OOD | ID | OOD | ID | OOD |
| Qwen2.5 (ZS) | 3.1 | 0.2 | 3.9 | 1.5 | 1.4 | 5.8 |
| Qwen2.5 (SFT) | 16.3 | 2.3 | 20.9 | 10.0 | 10.8 | 19.2 |
| Qwen2.5 (GRPO) | 22.5 | 2.0 | 30.6 | 7.5 | 20.0 | 26.8 |
| GPT-4.1 (ZS) | 13.0 | 3.4 | 17.3 | 5.0 | 8.0 | 21.5 |
| Gemini 2.5 Pro (ZS) | 28.1 | — | 36.0 | 13.5 | 24.9 | 33.7 |
| o4-mini (ZS) | 48.4 | — | 62.5 | 22.3 | 53.7 | 39.3 |
sum of EM and F1 to improve accuracy over time while still providing a useful training signal from difficult questions. The prompt for generating reasoning traces is provided in Appendix H. Since GRPO optimization is time-consuming, we restrict training to a single epoch which takes around 5 days.
The results in terms of exact match in Table 8 demonstrate that both SFT and GRPO slightly improve performance across dimensions, surpassing GPT-4.1 despite its much larger parameter size. However, Qwen2.5-7B-Instruct-1M is still not able to generalize to longer inputs, but does improve on questions related to out-of-distribution question topics and information processing skills.
5 Model Failure Analyses
We conduct additional analyses to understand model failures on SciTrek when using full-text articles as context. Table 9 shows correlations between model performance and various factors: the input article count, the question length (LenQ), the length of the underlying SQL query (LenSQL), and the answer length (LenA). They indicate that model performance correlates primarily with the question length, while o4-mini shows correlations with the input article count and the SQL query length.
| Model | #Articles | LenQ | LenSQL | LenA |
| Qwen2.5-7B-1M | 0.04 | 0.15∗ | 0.11 | 0.01 |
| Qwen2.5-14B-1M | 0.04 | 0.16∗ | 0.07 | 0.06 |
| Gemma-3-27B-IT | 0.08 | 0.14∗ | 0.02 | 0.08 |
| Llama-4-Scout | 0.10∗ | 0.17∗ | 0.03 | 0.03 |
| Llama-3.3-70B | 0.06 | 0.16∗ | 0.03 | 0.05 |
| DeepSeek-R1-Llama-70B | 0.15∗ | 0.16∗ | 0.02 | 0.08 |
| Gemini 2.5 Pro | 0.16∗ | 0.22∗ | 0.03 | 0.01 |
| GPT-4.1 | 0.07 | 0.21∗ | 0.06 | 0.06 |
| o4-mini | 0.14∗ | 0.09 | 0.25∗ | 0.06 |
| Qwen2.5-7B-1M (SFT) | 0.03 | 0.32∗ | 0.09 | 0.14∗ |
| Qwen2.5-7B-1M (GRPO) | 0.07 | 0.13∗ | 0.02 | 0.08 |
Inspection of model answers reveals several failure patterns: (1) weaker models tend to respond by simply outputting “NULL”, which suggests they rely more on the instruction to produce a fallback answer rather than genuinely understanding the given context; (2) models often fail to follow the specified output formats, e.g., returning author lists when author numbers are requested; (3) models sometimes provide incomplete answers when lists are requested for questions that require aggregation; and (4) models misinterpret compound conditions, particularly struggling with negation-based filtering (e.g., the last example in Table 1). Detailed analyses are provided in Appendix I.1.
We also manually analyzed the reasoning traces from our GRPO-based model (see examples in Appendix I.2). Specifically, we examined whether the reasoning is logically sound and individual steps obtain accurate information from specific contexts. We find that (1) the overall reasoning remains logically sound at 64K and 128K context lengths, although the model occasionally includes extraneous steps unrelated to the question; and (2) despite not being trained for 512K, the model still demonstrates coherent reasoning, but with reduced accuracy in specific steps compared to 64K and 128K; and (3) most reasoning steps that require counting are not accurate especially for references, which leads to incorrect answers. These findings indicate that GRPO-based reinforcement learning improves abstract reasoning, without enhancing the model’s accuracy in fine-grained operations, such as counting the number of references.
6 Conclusion
This paper introduced SciTrek, a benchmark designed for testing the ability of LLMs to perform multi-document information synthesis and structured reasoning over full-text scientific articles. By generating questions and ground-truth answers through a SQL backbone over article metadata, we provide a framework with explicit reasoning steps that is highly scalable and enables fine-grained error analysis. Our extensive evaluation demonstrates that SciTrek poses a significant challenge to both open-weight and proprietary LLMs, with only modest performance gains observed through supervised fine-tuning and reinforcement learning. Specifically, models struggle with compound logical conditions, fail at tasks requiring sorting, and often produce incomplete or badly formatted outputs. We believe the proposed methodology for creating SciTrek generalizes beyond titles, authors, and references to encompass broader metadata elements and domains, offering a tool to diagnose persistent shortcomings in LCLMs and improve their ability to support scientific workflows.
Appendix A Description of SciTrek’s Database
We construct a database with three tables for each article collection, using metadata from Semantic Scholar along with basic preprocessing. A description of these tables is provided in Table 10.
| Database Table | Column Name | Data Type | Description |
| articles | articleid | String | the unique identifier of the article |
| articletitle | String | the title of the article | |
| titlewordcount | Integer | the number of words in the article’s title (using spaces to determine word boundaries) | |
| authorcount | Integer | the number of authors in the article | |
| referencecount | Integer | the number of references that are cited in the article | |
| article-author | relationid | String | the unique identifier of the article-author relations |
| articleid | String | the identifier of the associated article | |
| authorname | String | the name of the author/s | |
| authorposition | Integer | the position of the author in the author list (starting from 0 for the first author) | |
| citing-cited | relationid | String | the unique identifier of the citation relations between two articles |
| articleidciting | String | the identifier of the article that cites another article | |
| articleidcited | String | the identifier of the article that is cited by another article |
Appendix B Prompts for Converting Between SQL Queries and Natural Language Questions
We generate natural language questions by prompting Qwen2.5-Coder-32B to convert SQL queries to natural language. Our prompt is presented in Figure 3. The prompt that we use to convert natural language questions back to to SQL queries is shown in Figure 4.
Appendix C Test Set Coverage of SQL Commands and Operators
Our test dataset covers all SQL commands and operators that are listed in Table 4. The detailed distribution is shown in Table 11.
| Command/Operator | #Instances | Proportion (%) |
| SELECT | 2,121 | 100.00% |
| WHERE | 1,821 | 85.86% |
| 1,032 | 48.66% | |
| IN | 682 | 32.15% |
| OR | 616 | 29.04% |
| ORDER BY | 591 | 27.86% |
| 476 | 22.44% | |
| 450 | 21.22% | |
| COUNT | 370 | 17.44% |
| ASC | 301 | 14.19% |
| DESC | 287 | 14.19% |
| DISTINCT | 280 | 13.20% |
| 215 | 10.14% | |
| MAX | 213 | 10.04% |
| AND | 195 | 9.19% |
| GROUP BY | 175 | 8.25% |
| 175 | 8.25% | |
| % | 164 | 7.73% |
| 154 | 7.26% | |
| NOT | 154 | 7.26% |
| 153 | 7.21% | |
| AVG | 135 | 6.36% |
| MIN | 133 | 6.27% |
| BETWEEN | 106 | 5.00% |
| SUM | 102 | 4.81% |
| 100 | 4.71% | |
| LIKE | 50 | 2.36% |
| + | 48 | 2.26% |
| - | 23 | 1.08% |
Appendix D Human Annotation Details
We validated the quality of our curated questions and answers with a human annotation study. We recruited crowdworkers via Prolific (https://www.prolific.com/), specifically native English speakers from the US or UK. Our annotators were compensated above the UK living wage, at £12 per hour. We randomly sampled 120 instances from our test dataset (20 per skill in Table 4). Each instance was annotated by three participants who were given the same instructions as those used for model testing with database tables as context (Figure 6, Appendix E). In each annotation session, crowdworkers were given three questions with different contexts. To ensure annotation quality, each session included a quality control question. On average, annotators spent about 5.5 minutes per session for the three questions.
Appendix E Prompts for Model Testing
We employ two different prompts for (zero-shot) model evaluation. The prompt for assessing LCLM capabilities against full-text articles is presented in Figure 5, and the prompt for using the corresponding database tables as the context is shown in Figure 6.
Appendix F Experimental Details
The details of the models we use in our experiments are presented in Table 12. For all models, we use their default inference configuration for answer generation from their Huggingface repositories.
We include both reasoning models and non-reasoning models. DeepSeek-R1-Distill-Llama-70B, Gemini 2.5 Pro and o4-mini are reasoning models. In our experiments, o4-mini conduct reasoning with the default thinking effort medium, Gemini 2.5 Pro was given a thinking budget of 512 tokens, and DeepSeek-R1-Distill-Llama-70B was prompted with the reasoning prompt in Figure 7.
| Model | #Parameters | Open Weights | Context Length | Release Date |
| Qwen2.5-7B-Instruct-1M | 7B | ✓ | 1,010,000 | Jan 2025 |
| Qwen2.5-14B-Instruct-1M | 14B | ✓ | 1,010,000 | Jan 2025 |
| Gemma-3-27B-IT | 27B | ✓ | 131,072 | Mar 2025 |
| Llama-4-Scout-17Bx16E-Instruct | 17Bx16E | ✓ | 10,485,760 | Apr 2025 |
| Llama-3.3-70B-Instruct | 70B | ✓ | 131,072 | Dec 2024 |
| DeepSeek-R1-Distill-Llama-70B | 70B | ✓ | 131,072 | Jan 2025 |
| Gemini 2.5 Pro | — | ✗ | 1,048,576 | Jun 2025 |
| GPT-4.1 | — | ✗ | 1,047,576 | Apr 2025 |
| o4-mini | — | ✗ | 200,000 | Apr 2025 |
Appendix G Additional Results
Table 13 presents additional results using F1 as the evaluation metric. Models are tested against full-text scientific articles and database tables as context. The input length averages only 1,980 tokens when database tables of are used as context across different lengths of full-text article collections.
| Full-text Articles | Database TablesΔ | ||||||||
| Model | CLen | 64K | 128K | 512K | 1M | 64K | 128K | 512K | 1M |
| Qwen2.5-7B-Instruct-1M | 1M | 8.7 | 7.0 | 1.1 | 0.0 | 39.2 | 39.3 | 28.7 | 21.2 |
| Qwen2.5-14B-Instruct-1M | 1M | 20.6 | 17.5 | 7.4 | 1.0 | 58.2 | 53.9 | 44.8 | 34.8 |
| Gemma-3-27B-IT | 128K | 22.6 | 14.7 | — | — | 52.8 | 52.5 | 47.8 | 40.7 |
| Llama-4-Scout-Instruct | 10M | 20.3 | 17.4 | 15.6 | 14.8 | 49.8 | 44.9 | 40.5 | 37.4 |
| Llama-3.3-70B-Instruct | 128K | 25.3 | 15.1 | — | — | 65.8 | 63.5 | 53.7 | 49.5 |
| DeepSeek-R1-Distill-Llama-70B | 128K | 29.0 | 9.0 | — | — | 85.5 | 77.6 | 68.0 | 62.3 |
| Gemini 2.5 Pro | 1M | 58.1 | 48.8 | 95.3 | 88.3 | 79.0 | 69.8 | ||
| GPT-4.1 | 1M | 36.0 | 29.7 | 22.3 | 19.6 | 82.8 | 73.3 | 61.8 | 56.3 |
| o4-mini | 195K | 74.0 | 63.8 | — | — | 97.2 | 89.2 | 87.0 | 86.5 |
Appendix H Reasoning Traces for Reinforcement Learning
The prompt used by GRPO to generate reasoning traces is given in Figure 7.
Appendix I Details for Model Failure Analyses
I.1 Analyses on Model Answers
To gain deeper insights into model behavior, we analyzed the failure patterns exhibited in model-generated answers on SciTrek. We randomly selected 60 instances in the input length of 128K that were hard for all zero-shot and supervised models (10 instances per skill in Table 4). We find that models tend to output “NULL” particularly on samples that require the skill of filtering. As shown in Table 14, most open-weight models tend to randomly output “NULL” for any sample. This indicates that these models rely more on the instruction to produce a fallback “NULL” answer rather than on genuine understanding of the provided context.141414Our testing instruction prompted models to generate “NULL” when they could not get the answer from the context. In addition, we find that models fail to follow the specified format for the answers, especially for sorting-related questions, as shown in Table 15. For example, models tend to generate author lists when sorted author numbers are requested and output lists when aggregates are requested. Interestingly, GPT-4.1 produces incorrectly formatted answers far more frequently than all other models. We further observe that some models tend to produce partial answers when lists are requested, especially for tasks that require aggregation. As shown in Table 16, our GRPO-based model tends to generate partial answers more than other models.
We also analyzed model performance on questions involving negation (e.g., those containing adverbs such as “not” or “never”). For questions that involve filtering-related skills, we have 132 instances with negation in our test set (Filtering: 11, Filtering+Aggregation: 15, Filtering+Sorting:12, and Relational Filtering:94). The results in Table 17 show that all models struggle with negation-based filtering.
| Aggregation | Sorting | Filtering | Filtering | Filtering | Relational | |
| Model | +Aggregation | +Sorting | Filtering | |||
| Qwen2.5-7B-Instruct-1M | 70 | 90 | 80 | 90 | 80 | 80 |
| Qwen2.5-14B-Instruct-1M | 10 | 10 | 20 | 50 | 50 | 60 |
| Gemma-3-27B-IT | 10 | 0 | 0 | 40 | 10 | 40 |
| Llama-4-Scout-14Bx16E-Instruct | 30 | 20 | 30 | 60 | 20 | 60 |
| Llama-3.3-70B-Instruct | 10 | 10 | 30 | 80 | 50 | 50 |
| DeepSeek-R1-Distill-Llama-70B | 0 | 0 | 0 | 10 | 0 | 10 |
| Gemini 2.5 Pro | 0 | 0 | 0 | 20 | 30 | 10 |
| GPT-4.1 | 0 | 0 | 0 | 10 | 0 | 0 |
| o4-mini | 0 | 0 | 10 | 10 | 30 | 20 |
| Qwen2.5-7B-Instruct-1M (SFT) | 0 | 0 | 40 | 30 | 40 | 0 |
| Qwen2.5-7B-Instruct-1M (GRPO) | 0 | 0 | 0 | 0 | 0 | 0 |
| Aggregation | Sorting | Filtering | Filtering | Filtering | Relational | |
| Model | +Aggregation | +Sorting | Filtering | |||
| Qwen2.5-7B-Instruct-1M | 0 | 0 | 0 | 0 | 0 | 0 |
| Qwen2.5-14B-Instruct-1M | 0 | 30 | 10 | 0 | 30 | 0 |
| Gemma-3-27B-IT | 0 | 20 | 30 | 20 | 20 | 0 |
| Llama-4-Scout-14Bx16E-Instruct | 10 | 10 | 0 | 10 | 30 | 10 |
| Llama-3.3-70B-Instruct | 10 | 10 | 30 | 0 | 10 | 0 |
| DeepSeek-R1-Distill-Llama-70B | 0 | 0 | 0 | 0 | 0 | 0 |
| Gemini 2.5 Pro | 0 | 10 | 0 | 0 | 0 | 0 |
| GPT-4.1 | 30 | 30 | 20 | 50 | 20 | 0 |
| o4-mini | 10 | 20 | 0 | 0 | 0 | 0 |
| Qwen2.5-7B-Instruct-1M (SFT) | 0 | 0 | 0 | 0 | 0 | 0 |
| Qwen2.5-7B-Instruct-1M (GRPO) | 0 | 0 | 0 | 10 | 0 | 0 |
| Aggregation | Sorting | Filtering | Filtering | Filtering | Relational | |
| Model | +Aggregation | +Sorting | Filtering | |||
| Qwen2.5-7B-Instruct-1M | 10 | 0 | 10 | 0 | 0 | 0 |
| Qwen2.5-14B-Instruct-1M | 0 | 0 | 0 | 0 | 0 | 10 |
| Gemma-3-27B-IT | 10 | 10 | 10 | 0 | 0 | 0 |
| Llama-4-Scout-14Bx16E-Instruct | 0 | 0 | 0 | 0 | 0 | 0 |
| Llama-3.3-70B-Instruct | 0 | 0 | 10 | 0 | 0 | 0 |
| DeepSeek-R1-Distill-Llama-70B | 0 | 0 | 0 | 20 | 0 | 0 |
| Gemini 2.5 Pro | 20 | 0 | 10 | 20 | 0 | 10 |
| GPT-4.1 | 10 | 0 | 0 | 0 | 0 | 0 |
| o4-mini | 10 | 0 | 20 | 30 | 10 | 0 |
| Qwen2.5-7B-Instruct-1M (SFT) | 10 | 0 | 0 | 0 | 0 | 20 |
| Qwen2.5-7B-Instruct-1M (GRPO) | 20 | 0 | 30 | 10 | 20 | 0 |
| Filtering | Filtering | Filtering | Relational | Average | |
| Model | +Aggregation | +Sorting | Filtering | ||
| Qwen2.5-7B-Instruct-1M | 0 | 0 | 0 | 0 | 0 |
| Qwen2.5-14B-Instruct-1M | 0 | 20.0 | 0 | 2.1 | 3.8 |
| Gemma-3-27B-IT | 0 | 6.7 | 0 | 0 | 0.8 |
| Llama-4-Scout-14Bx16E-Instruct | 0 | 20.0 | 0 | 2.1 | 3.8 |
| Llama-3.3-70B-Instruct | 0 | 6.7 | 0 | 0 | 0.8 |
| DeepSeek-R1-Distill-Llama-70B | 0 | 13.3 | 0 | 0 | 1.5 |
| Gemini 2.5 Pro | 0 | 20.0 | 8.3 | 6.4 | 9.1 |
| GPT-4.1 | 0 | 13.3 | 0 | 2.1 | 3.8 |
| o4-mini | 0 | 73.3 | 41.7 | 7.4 | 22.0 |
| Qwen2.5-7B-Instruct-1M (SFT) | 0 | 6.7 | 0 | 11.7 | 9.1 |
| Qwen2.5-7B-Instruct-1M (GRPO) | 0 | 33.3 | 8.3 | 3.2 | 8.3 |
I.2 Example Reasoning Traces from GRPO-based Model
We randomly sampled reasoning traces for three SQL query templates from the GRPO-based model. The reasoning traces for different sizes of contexts are presented in Table LABEL:tab:example_reasoning_traces.
[…]
Question
Context 64K |
What is the smallest number of authors for articles that have more than 13 references? |
[…]
To solve this problem, we need to follow these steps:
1. Identify articles with more than 13 references.
2. Determine the number of authors in each of those articles.
3. Find the smallest number of authors among these articles.
Let’s go through each step in detail:
### Step 1: Identify articles with more than 13 references
From the provided articles, we need to count the references in each article:
1. **”Best Constants in Young’s Inequality, Its Converse, and Its Generalization to More than Three Functions”**
- References: 13 (listed in the reference section)
2. **”Optical detection of strain and doping inhomogenieties in single layer MoS2”**
- References: 62 (listed in the reference section)
3. **”Ambipolar device simulation based on the drift-diffusion model in ion-gated transition metal dichalcogenide transistors”**
- References: 65 (listed in the reference section)
4. **”The Impact of AI in Physics Education: A Comprehensive Review from GCSE to University Levels”**
- References: 32 (listed in the reference section)
Only the articles titled ”Optical detection of strain and doping inhomogenieties in single layer MoS2”, ”Ambipolar device simulation based on the drift-diffusion model in ion-gated transition metal dichalcogenide transistors”, and ”The Impact of AI in Physics Education: A Comprehensive Review from GCSE to University Levels” have more than 13 references.
### Step 2: Determine the number of authors in each of these articles
Now, let’s count the authors in each of these articles:
1. **”Optical detection of strain and doping inhomogenieties in single layer MoS2”**
- Authors: A. Michail, N. Delikoukos, J. Parthenios, C. Galiotis, K. Papagelis (5 authors)
2. **”Ambipolar device simulation based on the drift-diffusion model in ion-gated transition metal dichalcogenide transistors”**
- Authors: Akiko Ueda, Yijin Zhang, Nobuyuki Sano, Hiroshi Imamura, Yoshihiro Iwasa (5 authors)
3. **”The Impact of AI in Physics Education: A Comprehensive Review from GCSE to University Levels”**
- Authors: Will Yeadon, Tom Hardy (2 authors)
### Step 3: Find the smallest number of authors
Among the articles with more than 13 references, the smallest number of authors is 2.
[…]
Question
Context 128K |
What is the smallest number of authors for articles that have more than 25 references? |
[…]
To solve this problem, we need to follow the steps outlined in the question. Let’s go through each step in detail.
### Step 1: Identify Articles with More Than 25 References
First, we need to go through each provided article and determine which ones have more than 25 references. We’ll count the words in the titles using spaces as delimiters to ensure accuracy.
#### Article Titles and Reference Counts:
1. **Article Title:** ”A SYSTEMATIC EVALUATION OF LARGE LANGUAGE MODELS OF CODE”
- Words: Count spaces to determine words.
A SYSTEMATIC EVALUATION OF LARGE LANGUAGE MODELS OF CODE
Spaces: 8 words (A, SYSTEMATIC, EVALUATION, OF, LARGE, LANGUAGE, MODELS, OF, CODE)
- Number of references: 27 (as given in the problem statement).
This article has more than 25 references.
2. **Article Title:** ”Reputation-based partner choice promotes cooperation in social networks”
- Words: Count spaces to determine words.
Reputation-based partner choice promotes cooperation in social networks
Spaces: 9 words (Reputation-based, partner, choice, promotes, cooperation, in, social, networks).
- Number of references: 26 (as given in the problem statement).
This article has more than 25 references.
<...we cut it down to save space...>
### Step 2: Determine the Smallest Number of Authors in Articles with More Than 25 References
Now that we’ve identified the articles with more than 25 references, let’s count the authors in each of these articles.
1. **Article Title:** ”A SYSTEMATIC EVALUATION OF LARGE LANGUAGE MODELS OF CODE”
- Authors: Frank F. Xu, Uri Alon, Graham Neubig, Vincent J. Hellendoorn (4 authors).
2. **Article Title:** ”Reputation-based partner choice promotes cooperation in social networks”
- Authors: Feng Fu, Christoph Hauert, Martin A. Nowak, Long Wang (4 authors).
3. **Article Title:** ”Efficient Multimodal Sampling via Tempered Distribution Flow”
- Authors: Yixuan Qiu, Xiao Wang (2 authors).
4. **Article Title:** ”Ultralow contact resistance between semimetal and monolayer semiconductors”
- Authors: Pin-Chun Shen, Cong Su, Yuxuan Lin, Ang-Sheng Chou, Chao-Ching Cheng, Ji-Hoon Park, Ming-Hui Chiu, Ang-Yu Lu, Hao-Ling Tang, Mohammad Mahdi Tavakoli, Gregory Pitner, Xiang Ji, Zhengyang Cai, Nannan Mao, Jiangtao Wang, Vincent Tung, Ju Li, Jeffrey Bokor, Alex Zettl, Chih-I Wu, Tomás Palacios, Lain-Jong Li, Jing Kong (11 authors).
5. **Article Title:** ”Normal forms for singularities of pedal curves produced by non-singular dual curve germs in Sn”
- Authors: Takashi Nishimura (1 author).
### Step 3: Identify the Smallest Number of Authors
[…]
### Final Answer:
[…]
Question
Context 512K |
What is the smallest number of authors for articles that cite more than 84 references? |
[…]
To solve this problem, we need to follow these steps:
1. Identify articles that cite more than 84 references.
2. Determine the number of authors in each qualifying article.
3. Identify the smallest number of authors among these articles.
4. Provide the final answer in the required format.
Let’s go through each step in detail.
### Step 1: Identify articles that cite more than 84 references
First, we will examine each provided article to determine which ones have more than 84 references.
1. **Article Title:** ”Macroscopic model and statistical model to characterize electromagnetic information of a digital coding metasurface”
Number of words in title: Counting spaces, we have ”Macroscopic model and statistical model to characterize electromagnetic information of a digital coding metasurface” → Words are: Macroscopic, model, and, statistical, model, to, characterize, electromagnetic, information, of, a, digital, coding, metasurface. (12 words).
Number of references: 85 (as stated in the problem).
2. **Article Title:** ”Worst-case VaR and robust portfolio optimization with interval random uncertainty set”
Number of words in title: ”Worst-case VaR and robust portfolio optimization with interval random uncertainty set” → Words are: Worst-case, VaR, and, robust, portfolio, optimization, with, interval, random, uncertainty, set. (11 words).
Number of references: 84 (as stated in the problem).
<...we cut it down to save space...>
Now we have identified the following articles that cite more than 84 references:
- ”Macroscopic model and statistical model to characterize electromagnetic information of a digital coding metasurface” (85 references).
- ”Worst-case VaR and robust portfolio optimization with interval random uncertainty set” (84 references).
- ”Direct Transmission of Digital Message via Programmable Coding Metasurface” (82 references).
- ”Bayesian experimental design for linear elasticity” (84 references).
- ”Metaverse for Social Good: A University Campus Prototype” (84 references).
- ”Alloying effect-induced electron polarization drives nitrate electroreduction to ammonia” (84 references).
### Step 2: Determine the number of authors in each qualifying article
Next, we will count the authors in each of these identified articles.
1. **”Macroscopic model and statistical model to characterize electromagnetic information of a digital coding metasurface”**
Authors listed: Rui Wen Shao, Wensheng Gan (listed as corresponding author), Philip S. Yu (listed as corresponding author).
Number of authors: 3.
2. **”Worst-case VaR and robust portfolio optimization with interval random uncertainty set”**
Authors listed: Wei Chen, Shaohua Tan, Dongqing Yang.
Number of authors: 3.
3. **”Direct Transmission of Digital Message via Programmable Coding Metasurface”**
Authors listed: Tie Jun Cui, Shuo Liu, Guo Dong Bai, Qian Ma.
Number of authors: 4.
4. **”Bayesian experimental design for linear elasticity”**
Authors listed: Haihan Duan, Jiaye Li, Sizheng Fan, Zhonghao Lin, Xiao Wu, Wei Cai (listed as corresponding author).
Number of authors: 6.
<...we cut it down to save space...>
### Step 3: Identify the smallest number of authors among these articles
Among the identified articles that cite more than 84 references, the smallest number of authors is 3.
### Final Answer
[…]
Question
Context 1M |
What is the smallest number of authors for articles that cite more than 51 references? |
[…]
Question
Context 64K |
How many different articles has Shi Jin been listed as the first author of? |
[…]
To determine how many different articles Shi Jin has been listed as the first author of, we need to go through each article provided in the documents and check his position as an author. Let’s examine each article one by:
1. **Broadband diffusion of terahertz waves by multi-bit coding metasurfaces**:
- Authors: Li-Hua Gao, Qiang Cheng, Jing Yang, Shao-Jie Ma, Jie Zhao, Shuo Liu, Hai-Bing Chen, Qiong He, Wei-Xiang Jiang, Hui-Feng Ma, Qi-Ye Wen, Lan-Ju Liang, Biao-Bing Jin, Wei-Wei Liu, Lei Zhou, Jian-Quan Yao, Pei-Heng Wu, Tie-Jun Cui
- Shi Jin is not listed as the first author here.
2. **Large Intelligent Surface-Assisted Wireless Communication Exploiting Statistical CSI**:
- Authors: Yu Han, Wankai Tang, Shi Jin, Chao-Kai Wen, Xiaoli Ma
- Shi Jin is listed as the fourth author but not the first.
3. **Intelligent Reflecting Surface Enhanced Wireless Network: Joint Active and Passive Beamforming Design**:
- Authors: Qingqing Wu, Rui Zhang
- Shi Jin is not listed as an author here.
4. **Wireless Communication Based on Information Metasurfaces**:
- Authors: Jun Yan Dai, Wankai Tang, Ming Zheng Chen, Chi Hou Chan, Qiang Cheng, Shi Jin, Tie Jun Cui
- Shi Jin is listed as the seventh author but not the first.
After examining all provided articles, Shi Jin has not been listed as the first author in any of them.
Thus, the number of different articles Shi Jin has been listed as the first author of is .
[…]
Question
Context 128K |
How many distinct articles has Tang been the first author of? |
[…]
To solve this problem, we need to follow a systematic approach by analyzing each article provided in the documents. Let’s go through each step in detail.
### Step 1: Identify Articles Written by Tang as the First Author
First, we need to examine each article to determine if Tang is listed as the first author. We will go through each article one by:
#### Article 1: ”The humping phenomenon during high speed gas metal arc welding”
Authors listed: T. C. Nguyen, D. C. Weckman, D. A. Johnson, H. W. Kerr (D. C. Weckman is the corresponding author).
Tang is not listed as an author in this article.
#### Article 2: ”Physics of Humping Formation in Laser Powder Bed Fusion”
Authors listed: C. Tang, K. Q. Le, C. H. Wong (C. Tang is the corresponding author).
Tang is listed as the first author in this article.
#### Article 3: ”Keyhole fluctuation and pore formation mechanisms during laser powder bed fusion additive manufacturing”
Authors listed: Yuze Huang 1,2, Tristan G. Fleming 3, Samuel J. Clark 1,2,4, Sebastian Marussi 1,2, Kamel Fezzaa 4, Jeyan Thiyagalingam 5, Chu Lun Alex Leung 1,2, Peter D. Lee 1,2 (Yuze Huang is the first author).
Tang is not listed as an author in this article.
#### Article 4: ”Unveiling mechanisms and onset threshold of humping in high-speed laser welding”
Authors listed: Zen-Hao Lai 1, Siguang Xu 2, Samuel J. Clark 3, Kamel Fezzaa 3, Jingjing Li 1,4 (Zen-Hao Lai is the first author).
Tang is not listed as an author in this article.
#### Article 5: ”Understanding of humping formation and suppression mechanisms using the numerical simulation”
Authors listed: Dongsheng Wu a,b,c, Xueming Hua a,b,∗, Dingjian Ye a,b, Fang Li a,b (Dongsheng Wu is the first author).
Tang is not listed as an author in this article.
#### Article 6: ”Investigation of the humping formation in the high power and high speed laser welding”
Authors listed: Yuewei Ai a,b, Ping Jia a,b,∗, Chunming Wang b, Gaoyang Mi b, Shaoning Geng a, Wei Liu a, Chu Han a (Yuewei Ai is the first author).
Tang is not listed as an author in this article.
### Step 2: Count Distinct Articles Written by Tang as the First Author
From the analysis above, only one article lists Tang as the first author:
- ”Physics of Humping Formation in Laser Powder Bed Fusion” by C. Tang, K. Q. Le, C. H. Wong.
Thus, Tang has been the first author in exactly one distinct article.
### Final Answer
[…]
Question
Context 512K |
How many different articles was Connor M. Brennan the first author of? |
[…]
To solve this problem, we need to follow the steps outlined in the question. Let’s go through each step carefully and determine the final answer.
### Step 1: Identify Relevant Articles
First, we need to scan through the provided articles to identify which ones have Connor M. Brennan as the first author. This will help us determine how many different articles he authored.
#### Article Titles and Authors:
1. **”The asymmetric effects of monetary policy on stock price bubbles”** by Christophe Blot, Paul Hubert, Fabien Labondance.
2. **”Deconstructing monetary policy surprises: the role of information shocks”** by Marek Jarociński, Peter Karadi.
3. **”Nominal Rigidities and the Dynamic Effects of a Shock to Monetary Policy”** by Lawrence J. Christiano, Martin Eichenbaum, Charles Evans.
4. **”Risk Appetite and the Risk-Taking Channel of Monetary Policy”** by Michael D. Bauer, Ben S. Bernanke, Eric Milstein.
5. **”The Macroeconomic Effects of the Federal Reserve’s Conventional and Unconventional Monetary Policies”** by Eric T. Swanson.
6. **”Federal Reserve Information and the Behavior of Interest Rates”** by Christina D. Romer, David H. Romer.
7. **”A Reassessment of Monetary Policy Surprises and High-Frequency Identification”** by Michael D. Bauer, Eric T. Swanson.
8. **”An Alternative Explanation for the ’Fed Information Effect’”** by Michael D. Bauer, Eric T. Swanson.
9. **”Monetary Policy Shocks: Data or Methods?”** by Connor M. Brennan, Margaret M. Jacobson, Christian Matthes, Todd B. Walker.
10. **”High Asymmetric Effects of Monetary Policy on Stock Price Bubbles”** by Christophe Blot, Paul Hubert, Fabien Labondance (Note: This is a duplicate title and should be ignored).
From this list, only one article has Connor M. Brennan as the first author:
- **”Monetary Policy Shocks: Data or Methods?”** by Connor M. Brennan, Margaret M. Jacobson, Christian Matthes, Todd B. Walker.
### Step 2: Verify Connor M. Brennan as the First Author
Now, let’s confirm that Connor M. Brennan is indeed the first author of the identified article:
- The article title is ”Monetary Policy Shocks: Data or Methods?”
- The authors listed are Connor M. Brennan, Margaret M. Jacobson, Christian Matthes, Todd B. Walker.
- Connor M. Brennan is listed first, confirming he is the first author.
### Step 3: Count Words in the Title Using Spaces
Next, we need to count the words in the title ”Monetary Policy Shocks: Data or Methods?” using spaces as delimiters. Let’s list each word:
1. Monetary
2. Policy
3. Shocks:
4. Data
5. or
6. Methods?
There are 6 words in the title.
### Final Answer
[…]
Question
Context 1M |
How many different articles has Saurabh Bodhe authored as the first author? |
[…]
Question
Context 64K |
How many different articles has Feng Miao authored? |
[…]
To determine how many different articles Feng Miao has authored, we need to go through each article provided and identify Feng Miao’s involvement in each one. Let’s analyze each article one by step.
### Article 1: ”Hopping transport through defect-induced localized states in molybdenum disulphide”
Authors listed: Hao Qiu, Tao Xu, Zilu Wang, Wei Ren, Haiyan Nan, Zhenhua Ni, Qian Chen, Shijun Yuan, Feng Miao, Fengqi Song, Gen Long, Yi Shi, Litao Sun, Jinlan Wang, Xinran Wang.
Feng Miao is listed as an author in this article.
### Article 2: ”Approaching the quantum limit in two-dimensional semiconductor contacts”
Authors listed: Weisheng Li, Xiaoshu Gong, Zhihao Yu, Liang Ma, Wenjie Sun, Si Gao, Çağıl Köroğlu, Wenfeng Wang, Lei Liu, Taotao Li, Hongkai Ning, Dongxu Fan, Yifei Xu, Xuecou Tu, Tao Xu, Litao Sun, Wenhui Wang, Junpeng Lu, Zhenhua Ni, Jia Li, Xidong Duan, Peng Wang, Yuefeng Nie, Hao Qiu, Yi Shi, Eric Pop, Jinlan Wang, Xinran Wang.
Feng Miao is not listed as an author in this article.
### Article 3: ”Origins of genuine Ohmic van der Waals contact between indium and MoS2”
Authors listed: Bum-Kyu Kim, Tae-Hyung Kim, Dong-Hwan Choi, Hanul Kim, Kenji Watanabe, Takashi Taniguchi, Heesuk Rho, Ju-Jin Kim, Yong-Hoon Kim, Myung-Ho Bae.
Feng Miao is not listed as an author in this article.
### Article 4: ”Van der Waals contacts between three-dimensional metals and two-dimensional semiconductors”
Authors listed: Yan Wang, Jong Chan Kim, Ryan J. Wu, Jenny Martinez, Xiuju Song, Jieun Yang, Fang Zhao, Andre Mkhoyan, Hu Young Jeong, Manish Chhowalla.
Feng Miao is not listed as an author in this article.
### Conclusion
Based on the analysis, Feng Miao has authored only one article, which is ”Hopping transport through defect-induced localized states in molybdenum disulphide.”
[…]
Question
Context 128K |
How many different articles has JIAYI SUN authored? |
[…]
To solve this problem, we need to follow these steps:
1. Identify articles authored by JIAYI SUN.
2. Count the number of distinct articles authored by JIAYI SUN.
3. Provide the final answer in the required format.
Let’s go through each step in detail:
### Step 1: Identify articles authored by JIAYI SUN
From the provided articles, we need to examine each one to determine if JIAYI SUN is listed as an author. Let’s review each article:
1. **Article Title:** ”Solipsis: A Decentralized Architecture for Virtual Environments”
Authors listed: Davide Frey, Jérôme Royan, Romain Piegay, Anne-Marie Kermarrec, Emmanuelle Anceaume, Fabrice Le Fessant, and Jiayi Sun.
JIAYI SUN is present here.
2. **Article Title:** ”Metaverse: Survey, Applications, Security, and Opportunities”
Authors listed: Jiayi Sun, Wensheng Gan, Han-Chieh Chao, and Philip S. Yu.
JIAYI SUN is present here.
3. **Article Title:** ”Big Data Meets Metaverse: A Survey”
Authors listed: Jiayi Sun, Wensheng Gan, Zefeng Chen, Junhui Li, and Philip S. Yu.
JIAYI SUN is present here.
4. **Article Title:** ”A Dynamic Resource Allocation Framework for Synchronizing Metaverse with IoT Service and Data”
Authors listed: Yue Han, Dusit Niyato, Cyril Leung, Chunyan Miao, and Jiayi Sun.
JIAYI SUN is present here.
Thus, JIAYI SUN has authored four distinct articles.
### Step 2: Count the number of distinct articles authored by JIAYI SUN
From the analysis above, we have identified the following articles authored by JIAYI SUN:
1. ”Solipsis: A Decentralized Architecture for Virtual Environments”
2. ”Metaverse: Survey, Applications, Security, and Opportunities”
3. ”Big Data Meets Metaverse: A Survey”
4. ”A Dynamic Resource Allocation Framework for Synchronizing Metaverse with IoT Service and Data”
[…]
### Final Answer
[…]
Question
Context 512K |
How many different articles has Qiang Cheng authored? |
[…]
To solve this problem, we need to follow a systematic approach by analyzing each provided article to identify instances where Qiang Cheng is listed as an author. Let’s go through each step in detail.
### Step 1: Identify Articles Mentions Qiang Cheng as an Author
First, we will examine each article title and author list to determine which articles include Qiang Cheng as an author.
1. **Article Title:** ”Intelligent Reflecting Surface Enhanced Wireless Network: Joint Active and Passive Beamforming Design”
[…]
2. **Article Title:** ”Broadband continuous beam-steering with time-modulated metasurfaces in the near-infrared spectral regime”
[…]
3. **Article Title:** ”Path Loss Modeling and Measurements for Reconfigurable Intelligent Surfaces in the Millimeter-Wave Frequency Band”
[…]
4. **Article Title:** ”Index Modulation for 6G Wireless Networks Using Time-Varying Metasurfaces”
[…]
From this analysis, we see that Qiang Cheng is an author in four different articles.
### Step 2: Count the Number of Articles Featuring Qiang Cheng
Now that we have identified the articles where Qiang Cheng is an author, we can count them. The list of articles includes:
1. ”Intelligent Reflecting Surface Enhanced Wireless Network: Joint Active and Passive Beamforming Design”
2. ”Broadband continuous beam-steering with time-modulated metasurfaces in the near-infrared spectral regime”
3. ”Path Loss Modeling and Measurements for Reconfigurable Intelligent Surfaces in the Millimeter-Wave Frequency Band”
4. ”Index Modulation for 6G Wireless Networks Using Time-Varying Metasurfaces”
[…]
### Final Answer
[…]
Question
Context 1M |
How many different articles has A. N. Pettitt contributed to? |
[…]

### v2

@@
-Who Gets Cited Most? Benchmarking Long-Context Language Models on Scientific Articles
+Who Gets Cited Most? Benchmarking Long-Context Reasoning on Scientific Articles
 Abstract
-This paper introduces SciTrek, a novel question-answering benchmark designed to evaluate the long-context reasoning capabilities of large language models (LLMs) using scientific articles.
-Current long-context benchmarks often rely on non-scientific texts, focus on simple information retrieval tasks, or employ artificial contexts.
-SciTrek addresses these limitations by proposing complex questions that require information aggregation and synthesis across multiple full-text scientific articles.
-Questions and their ground-truth answers are automatically generated by formulating them as SQL queries over a database constructed from article metadata (titles, authors, and references).
-The SQL operations provide explicit, verifiable reasoning steps for fine-grained error analysis, and the construction process scales to contexts up to 1M tokens with minimal supervision.
-Extensive experiments on a diverse set of open-weight and proprietary LLMs demonstrate that SciTrek poses a significant challenge as the context length increases, with supervised fine-tuning and reinforcement learning offering only limited gains.
-Our analysis reveals systematic shortcomings in models’ abilities to perform basic numerical operations and accurately locate specific information in long contexts.111Our dataset can be downloaded from https://github.com/oaimli/SciTrek
-1 Introduction
-Large language models (LLMs) show promise in accelerating scientific progress by assisting researchers in reviewing and synthesizing the growing body of literature (Pearson, 2024; Agarwal et al., 2025) and generating novel research ideas (Si et al., 2025; Baek et al., 2025).
-So much so that industry products such as Deep Research222https://openai.com/index/introducing-deep-research/, Elicit333https://elicit.com/, and Scite444https://scite.ai/ have emerged as dedicated tools to aid users with complex, multi-step research tasks.
-Scientific workflows often require processing large inputs: full scientific articles, collections of references, or structured datasets.
-A model capable of processing 100K+ tokens could simultaneously analyze multiple articles, track long chains of reasoning, or connect experimental results to prior work.
-While long-context language models (LCLMs) are highly relevant to scientific tasks, existing benchmarks designed to evaluate their capabilities predominantly focus on non-scientific texts (Maekawa et al., 2025; Bai et al., 2025; Yen et al., 2025; Kuratov et al., 2024).
-They mostly address simple information retrieval tasks (e.g., Needle-In-A-Haystack; Kamradt 2023; Hsieh et al.
-2024) rather than information integration, i.e., reasoning across multiple documents to synthesize information and answer complex questions.
-Some question-answering benchmarks (Asai et al., 2024; Bai et al., 2025; An et al., 2024) require information integration, but do not provide a typology of the reasoning steps involved, which makes it difficult to determine the specific capabilities being assessed and why a model fails.
-Another issue relates to scalability and extensibility.
-When substantial human effort is involved in generating the questions and their answers (e.g., Asai et al.
-2024), scaling to longer contexts or different types of questions becomes prohibitive.
-To better emulate scientific research practice, it is also important to evaluate a range of skills in natural contexts, such as a model’s ability to aggregate or filter information or combinations thereof (Wang et al., 2024c).
-However, a few benchmarks rely on artificial contexts, often created by extending short-context tasks (e.g., HotpotQA; Yang et al.
-2018) with added noise (Kuratov et al., 2024) or by synthesizing the input with LLMs (Maekawa et al., 2025).
-In this paper, we create a question-answering benchmark based on scientific articles that aims to alleviate these issues.
-We assume a scenario where a researcher, possibly during the literature review process, seeks answers to specific questions about the scientific articles they read.
-As an example, consider the first question in Table 1: “What is the highest number of authors that any single article has?”.
-Although seemingly simple, it requires aggregation of information across multiple articles.
-To answer correctly, a hypothetical model would need to count the number of authors for each article within a document collection, and identify the article with the largest number.
-To generate questions with natural language inputs and explicit reasoning steps, we focus on basic numerical operations such as counting, aggregating, and sorting, applied to key elements of scientific articles: titles, authors, and references.
-Since these operations correspond to standard database functions, we can automatically generate arbitrarily complex SQL queries over any collection of scientific articles and obtain ground-truth answers without manual annotation by constructing database tables for titles, authors, and references.
-We take advantage of recent advances in LLM capabilities for SQL understanding and generation (Yu et al., 2018; Li et al., 2023; Hui et al., 2024) to automatically convert SQL queries into corresponding natural language questions.
+We introduce SciTrek, a novel question-answering benchmark designed to evaluate long-context reasoning capabilities of large language models (LLMs) using scientific articles.
+Current long-context benchmarks often focus on simple information retrieval tasks, or employ artificial contexts.
+SciTrek addresses these limitations by creating benchmark questions that require information aggregation and synthesis across multiple full-text scientific articles.
+The questions and their ground-truth answers are automatically generated by formulating them as SQL queries over a database constructed from article metadata (i.e., titles, authors, and references).
+These SQL queries provide explicit, verifiable reasoning processes that enable fine-grained error analysis on model answers, and the data construction scales to contexts of up to 1M tokens with minimal supervision.
+Experiments on open-weight and proprietary LLMs show that SciTrek poses significant challenges as the context length increases, with supervised fine-tuning and reinforcement learning offering only limited gains.
+Our analysis reveals systematic shortcomings of frontier LLMs’ ability to effectively perform numerical operations and accurately locate information in long contexts.111Our dataset and code are available from https://github.com/oaimli/SciTrek
 | Question | SQL Query | Answer |
@@
 | What is the word count of the titles of articles, sorted by the number of authors in ascending order? | SELECT title_word_count FROM articles ORDER BY author_count ASC | 9, 17, 5, 6, 9, 12 |
-| How many references do articles with exactly two authors have? | SELECT reference_count FROM articles WHERE author_count = 2 | 16 |
+| How many references do articles with exactly two authors have? | SELECT reference_count FROM articles WHERE author_count = 2 |
+16
 | What is the total number of words in the titles of all articles that have exactly 60 references? | SELECT SUM(title_word_count) FROM articles WHERE reference_count 60 |
@@
 Vaswani |
-| How many articles have been cited by other articles but do not cite any other articles?
+| How many articles are cited by other articles but do not cite any other articles?
 | SELECT COUNT(*) FROM articles WHERE article_id NOT IN (SELECT article_id_citing FROM citing_cited) AND article_id IN (SELECT article_id_cited FROM citing_cited) |
 2
-Our benchmark, SciTrek, contains a test set of 2,121 question-answer pairs on scientific articles and a training set of 19,543 question-answer pairs, with contexts of varying lengths (i.e., 64K, 128K, 512K, and 1M555Throughout the paper, K refers to 1,024 and M to 1,024K.).
-Table 1 illustrates examples of questions, their corresponding SQL queries, and answers.
-These questions are relatively superficial, as they do not engage deeply with article content.
-However, the metadata consists of distillable facts sparsely distributed across article collections, requiring models to retrieve and synthesize information over long contexts.
-If models struggle with simple filtering and sorting, they are unlikely to succeed on more complex analytical tasks.
-We use SciTrek to evaluate the capabilities of LCLMs with context windows exceeding 128K tokens.
-Experimental results reveal that the benchmark poses significant challenges, with models achieving poor performance, even when enhanced with supervised fine-tuning (SFT) and reinforcement learning (RL).
-We make the following contributions:
-We propose a methodology for constructing a long-context question-answering benchmark over natural texts, with explicit reasoning steps approximated via SQL operations that involve information aggregation.
-The methodology requires minimal human intervention and can be replicated for other domains containing entities that can be structured as database tables.
-Extensive experiments reveal that frontier open-weight and proprietary LLMs struggle significantly with this task, especially as input length increases.
+1 Introduction
+Large language models (LLMs) show promise in accelerating scientific progress by assisting researchers in reviewing and synthesizing the growing body of literature (Pearson, 2024; Agarwal et al., 2025) and generating novel research ideas (Si et al., 2025; Baek et al., 2025).
+So much so that industry products such as Deep Research222https://openai.com/index/introducing-deep-research/, Elicit333https://elicit.com/, and Scite444https://scite.ai/ have emerged as dedicated tools to aid users with complex, multi-step research tasks.
+Scientific workflows often require processing large inputs: full scientific articles, collections of references, or structured datasets.
+A model capable of processing 100K+ tokens could simultaneously analyze multiple articles, track long chains of reasoning, or connect experimental results to prior work.
+While long-context language models (LCLMs) are highly relevant to scientific tasks, existing benchmarks for evaluating their capabilities predominantly focus on non-scientific texts (Kuratov et al., 2024; Maekawa et al., 2025; Bai et al., 2025; Yen et al., 2025).
+They mostly address simple information retrieval tasks (e.g., Needle-In-A-Haystack; Kamradt 2023; Hsieh et al.
+2024) rather than information integration, i.e., reasoning across multiple documents to synthesize information and answer complex questions.
+Some question-answering benchmarks (Asai et al., 2024; An et al., 2024; Bai et al., 2025) require information integration but do not provide a typology of the reasoning steps involved, making it difficult to determine which capabilities are being assessed and why a model fails.
+Another issue relates to scalability and extensibility.
+When substantial human effort is required to generate benchmark questions and their answers (e.g., Asai et al.
+2024), scaling to longer contexts or different types of questions becomes prohibitive.
+To better emulate scientific research practice, it is also important to evaluate a range of skills in natural contexts, such as a model’s ability to aggregate or filter information, and combinations thereof (Wang et al., 2024c).
+In contrast, a few benchmarks rely on artificial contexts, created by adding noise (Kuratov et al., 2024) to short contexts (e.g., HotpotQA; Yang et al.
+2018) or by synthesizing the input with LLMs (Maekawa et al., 2025).
+In this paper, we create a question-answering benchmark based on scientific articles that aims to alleviate these issues.
+We assume a scenario where a researcher, possibly during the literature review process, seeks answers to specific questions when the related scientific articles are provided.
+As an example, consider the first question in Table 1: “What is the highest number of authors that any single article has?”.
+Although seemingly simple, it requires aggregation of information across multiple articles.
+To answer correctly, a hypothetical model would need to count the authors of each article in the collection and select the one with the largest number.
+We generate questions requiring explicit reasoning by focusing on basic numerical operations such as counting, aggregating, and sorting, applied to core elements of scientific articles: titles, authors, and references.
+Since these operations correspond to standard database functions, we can automatically generate arbitrarily complex SQL queries over any collection of scientific articles and obtain ground-truth answers without manual annotation by constructing database tables for titles, authors, and references.
+We exploit recent advances in LLM capabilities for SQL understanding and generation (Yu et al., 2018; Li et al., 2023; Hui et al., 2024) to automatically convert SQL queries into natural-language questions.
+Our benchmark, SciTrek, contains a test set of 2,121 question-answer pairs on scientific articles and a training set of 19,543 question-answer pairs, with contexts of varying lengths (i.e., 64K, 128K, 512K, and 1M).555In the paper, K refers to 1,024 and M to 1,024K.
+Table 1 shows examples of questions, their corresponding SQL queries, and answers.
+While these questions are relatively superficial and do not engage deeply with the domain-specific content, they require models to retrieve and synthesize facts that are sparsely distributed across article collections and span long contexts.
+If models struggle with simple comparison, filtering, sorting, and aggregation operations, they might be impossible to succeed on more complex analytical tasks (Wolfson et al., 2020, 2025).
+We use SciTrek to evaluate the reasoning capabilities of LCLMs with context windows exceeding 128K tokens.
+Experimental results reveal that SciTrek poses significant challenges for these LCLMs.
+In general, models achieve poor performance, even when enhanced with supervised fine-tuning (SFT) and reinforcement learning (RL).
+This paper makes the following contributions:
+We propose a methodology for constructing a long-context question-answering benchmark over natural texts, where explicit reasoning processes are approximated via SQL operations involving information aggregation.
+The methodology requires minimal human intervention and generalizes to any domain with entities that can be structured as database tables.
+Extensive experiments reveal that frontier open-weight and proprietary LLMs struggle with this task, especially as input length increases.
 While SFT and RL improve over baseline models, performance remains limited.
-All models struggle with questions involving references and citation relationships among the input articles.
+All models perform poorly on questions involving references and citation relationships among articles.
 By design, SciTrek enables fine-grained analysis of model behavior, offering insight into where and why models fail.
-Our evaluation of open-weight models reveals systematic shortcomings in counting bibliographic elements and performing basic numerical operations, with error rates increasing as question length grows.
-Interestingly, models frequently misinterpret compound conditions and struggle with logical constructs involving negation.
-(see the last question in Table 1).
-| Dataset | Natural | Skills | Reason | Scalable | Length | Example Question |
-| NeedleBench | ✗ | ✗ | ✗ | ✓ | 128K | What legendary item is hidden on Emerald Island? |
-| Ada-LEval | ✗ | ✗ | ✗ | ✓ | 128K | What is the correct order of the segments? |
-| BABILong | ✗ | ✗ | ✗ | ✓ | 10M | Where is Mary? |
-| HELMET | ✗ | ✓ | ✗ | ✗ | 128K | Who set the fire in one tree hill? |
-| LIFBENCH | ✗ | ✓ | ✗ | ✓ | 128K | Retrieve the entry at position 8th in the list. |
-| RULER | ✗ | ✓ | ✗ | ✓ | 128K | Find all variables that are assigned the value 12345. |
-| OpenScholar | ✓ | ✓ | ✗ | ✗ | 3K∗ | Compile a list of reviews […], and identify the most promising […] |
-| LongBench v2 | ✓ | ✓ | ✗ | ✗ | 128K | How long have I been living in my current apartment in Shinjuku? |
-| LongMemEval | ✓ | ✓ | ✗ | ✗ | 2M | How many bikes do I currently own? |
-| L-Eval | ✓ | ✓ | ✗ | ✗ | 200K | How do I know when I should apply for Medicare? |
-| HoloBench | ✗ | ✓ | ✓ | ✓ | 64K | What are the names of wines and their corresponding grape types? |
-| MathHay | ✗ | ✓ | ✓ | ✓ | 128K | What is the total number of points scored by LeBron […] combined? |
-| Loong | ✓ | ✗ | ✓ | ✓ | 250K | Which company has the highest non-current assets? |
-| SciTrek | ✓ | ✓ | ✓ | ✓ | 1M | How many articles have been cited by other articles but do not cite any other articles? |
+Evaluation of open-weight models reveals systematic shortcomings in counting bibliographic elements and performing basic numeric operations, with error rates increasing as question length grows; models also frequently misinterpret compound conditions and struggle with logical constructs involving negation (see last question in Table 1).
+| Dataset | Natural Context | #SciQs | Reasoning | Scalable | Len | Example Question |
+| NeedleBench | ✗ | — | ✗ | ✓ | 128K | What legendary item is hidden on Emerald Island? |
+| Ada-LEval | ✗ | — | ✗ | ✓ | 128K | What is the correct order of the segments? |
+| BABILong | ✗ | — | ✗ | ✓ | 10M | Where is Mary? |
+| HELMET | ✗ | — | ✗ | ✗ | 128K | Who set the fire in one tree hill? |
+| LIFBENCH | ✗ | — | ✗ | ✓ | 128K | Retrieve the entry at position 8th in the list. |
+| RULER | ✗ | — | ✗ | ✓ | 128K | Find all variables that are assigned the value 12345. |
+| OpenScholar | ✓ | 208 | ✗ | ✗ | 3K∗ | Compile a list of reviews […], and identify the most promising […] |
+| LongBench v2 | ✓ | 50 | ✗ | ✗ | 128K | How long have I been living in my current apartment in Shinjuku? |
+| LongMemEval | ✓ | — | ✗ | ✗ | 2M | How many bikes do I currently own? |
+| L-Eval | ✓ | — | ✗ | ✗ | 200K | How do I know when I should apply for Medicare? |
+| HoloBench | ✗ | — | ✓ | ✓ | 64K | What are the names of wines and their corresponding grape types? |
+| MathHay | ✗ | — | ✓ | ✓ | 128K | What is the total number of points scored by LeBron […] combined? |
+| Loong | ✓ | 53 | ✓ | ✓ | 250K | Which company has the highest non-current assets? |
+| SciTrek | ✓ | 1,716 | ✓ | ✓ | 1M | How many articles are cited by other articles but do not cite any other articles? |
 2 Related Work
@@
 Rotary Position Embeddings (RoPE; Su et al.
-2024) have emerged as a widely used approach, replacing absolute positional embeddings with rotational transformations of token embeddings.
-It improves generalization to unseen sequence lengths.
+2024) have emerged as a widely used approach that replaces absolute positional embeddings with rotational transformations of token embeddings, improving generalization to unseen sequence lengths.
 Building on RoPE, several extensions introduce rescaling to further enhance performance on substantially longer inputs without the need for retraining (Peng et al., 2024).
 Other architectural advances focus on reducing the quadratic cost of the Transformer attention mechanism (Sun et al., 2025).
-Sparse attention methods, such as LongFormer (Beltagy et al., 2020; Jiang et al., 2024; Fu et al., 2024), compute only selected portions of the full attention matrix, thereby reducing computational cost and accelerating both prefilling and inference.
-Many contemporary models interleave sparse and full attention layers to balance efficiency and performance (Llama Team, 2025; Yang et al., 2025; Kamath et al., 2025; Dubey et al., 2024).
-Another popular strategy of reducing the computation required in long-context settings is to limit the number of active model parameters.
+Sparse attention methods, such as LongFormer (Beltagy et al., 2020), compute only selected portions of the full attention matrix, thereby reducing computational cost and accelerating both prefilling and inference (Jiang et al., 2024; Fu et al., 2024).
+Many contemporary models interleave sparse and full attention layers to balance efficiency and performance (Dubey et al., 2024; Llama Team, 2025; Yang et al., 2025; Kamath et al., 2025).
+Another popular strategy for reducing computation in long-context settings is to limit the number of active model parameters.
 Group Query Attention (GQA; Ainslie et al.
 2023) enables multiple attention heads to share key–value projections, without significant performance loss.
-The open-weight LLMs used in our experiments adopt combinations of these techniques, complemented by pretraining on long-context tasks (Dubey et al., 2024; Kamath et al., 2025; Yang et al., 2025; Llama Team, 2025).
+The open-weight LLMs used in our experiments combine these techniques, complemented by pretraining on long-context tasks (Dubey et al., 2024; Kamath et al., 2025; Yang et al., 2025; Llama Team, 2025).
 Long-context Datasets
-With growing interest in LCLMs, there has been a surge in benchmarks designed to evaluate their performance.
-As far as text-based efforts are concerned, initial benchmarks primarily assessed whether models are able to retrieve relevant information from their context, typically through Needle-In-A-Haystack (NIAH) tasks, such as NeedleBench (Li et al., 2024).
-Building on these, subsequent benchmarks introduced either distractor information to short-context tasks or dispersed relevant information across synthetic contexts, such as Ada-LEval (Wang et al., 2024a), BABILong (Kuratov et al., 2024), HELMET (Yen et al., 2025), LIFBENCH (Wu et al., 2025b), LongBench (Bai et al., 2024), and RULER (Hsieh et al., 2024).
+The growing interest in LCLMs has driven the development of numerous benchmarks designed to evaluate their performance.
+Initial benchmarks primarily assessed whether models can retrieve information from their context, typically through Needle-In-A-Haystack (NIAH) tasks, such as NeedleBench (Li et al., 2024).
+Building on these approaches, subsequent benchmarks either added distractor information to short-context tasks or scattered relevant information across synthetic contexts, such as Ada-LEval (Wang et al., 2024a), BABILong (Kuratov et al., 2024), HELMET (Yen et al., 2025), LIFBENCH (Wu et al., 2025b), LongBench (Bai et al., 2024), and RULER (Hsieh et al., 2024).
 In addition to retrieval, some benchmarks emphasize complex reasoning over long contexts.
-For example, OpenScholar (Asai et al., 2024) and LongBench v2 (Bai et al., 2025) contains questions that require extensive use of the multiple documents provided as context, while DocFinQA (Reddy et al., 2024) and MedOdyssey (Fan et al., 2025) focus on financial and medical reasoning, respectively.
-Although not explicitly technical, LongMemEval (Wu et al., 2025a) and L-Eval (An et al., 2024) also requires reasoning over natural, long-context information.
-However, evaluating model answers against these benchmarks remains challenging, or the reasoning steps required to arrive at correct responses are often opaque, making it difficult to assess failure modes of LCLMs.
+For example, OpenScholar (Asai et al., 2024) and LongBench v2 (Bai et al., 2025) contain questions that require extensive use of multiple documents provided as context, while DocFinQA (Reddy et al., 2024) and MedOdyssey (Fan et al., 2025) focus on financial and medical reasoning, respectively.
+Although not explicitly technical, LongMemEval (Wu et al., 2025a) and L-Eval (An et al., 2024) also require reasoning over natural, long-context information.
+However, evaluating model answers against these benchmarks remains challenging, as the reasoning steps required to arrive at correct responses are often opaque, making it difficult to assess failure modes.
 In contrast, HoloBench (Maekawa et al., 2025) and MathHay (Wang et al., 2024b) better delineate between the types of reasoning required for each question, but rely on unnatural contexts.
-Loong (Wang et al., 2024c) combines structured reasoning and natural contexts, but tests a very limited number of skills.
-CURIE (Cui et al., 2025), OpenScholar (Asai et al., 2024) and LongBench v2 (Bai et al., 2025) evaluate reasoning over scientific articles, but their reliance on expert-curated questions and answers poses challenges for scalability to longer contexts or larger datasets.
+CURIE (Cui et al., 2025), OpenScholar, and LongBench v2 evaluate reasoning over scientific articles, but their reliance on expert-curated questions and answers poses challenges for scalability to longer contexts or larger datasets.
 In designing SciTrek, we aim to retain many of the useful qualities of previous benchmarks (e.g., structured reasoning, scalability) while still proposing a fairly natural and easy task for humans to perform.
-Although we focus on scientific articles, our methodology extends to other domains with explicit entities and relations.
-Thanks to the SQL backbone, we are able to construct questions testing various model skills (e.g., aggregation, filtering), and how these manifest themselves through explicit reasoning steps.
-We compare SciTrek with representative existing benchmarks in Table 2.
+We compare SciTrek with existing benchmarks in Table 2.
+Unlike Loong (Wang et al., 2024c), which tests broad long-context comprehension, SciTrek uniquely targets scientific synthesis and structured reasoning at extreme context lengths.
+Although we focus on scientific articles, our methodology could extend to other domains with explicit entities and relations.
+Thanks to the SQL backbone, we are able to construct questions testing various model skills (e.g., aggregation, filtering), and how these manifest themselves through explicit reasoning processes.
 3 The SciTrek Benchmark
-In this section, we explain how SciTrek was curated.
-As mentioned earlier, our benchmark consists of question-answer pairs over scientific articles.
-Figure 1 illustrates our process: we first gather scientific article collections as contexts corresponding to different lengths (i.e., 64K, 128K, 512K, and 1M); we then create databases representing the article metadata and SQL queries with answers from database execution; finally, we convert the SQL queries into natural language questions.
+Figure 1 illustrates how SciTrek was curated: we first gather scientific article collections as contexts corresponding to different lengths (i.e., 64K, 128K, 512K, and 1M); we then create databases representing the article metadata and SQL queries with answers from database execution; finally, we convert the SQL queries into natural language questions.
+We explain these stages in detail below.
 3.1 Gathering Scientific Article Collections
-SciTrek is constructed using scientific articles from Semantic Scholar666https://www.semanticscholar.org/.
-To cover diverse topics of articles, we obtain an initial set of seed articles from eight subjects: Computer Science (CS), Economics, Electronic Engineering (EE), Math, Physics, Biology, Finance, and Statistics.777Following the typology of subjects from https://arxiv.org/.
+SciTrek is constructed using scientific articles from Semantic Scholar.666https://www.semanticscholar.org/ To cover diverse topics of articles, we obtain an initial set of seed articles from eight subjects: Computer Science (CS), Economics, Electronic Engineering (EE), Math, Physics, Biology, Finance, and Statistics.777Following the typology of subjects from https://arxiv.org/.
 For each subject, we select two seed articles with more than 100 citations since 2020.
@@
 Finally, we convert the collected PDFs into markdown texts using Marker.999https://github.com/datalab-to/marker
-From these clusters, we construct article collections of varying context lengths for our question-answering task.
+From these clusters, we create article collections of varying context lengths for our question-answering task.
 By concatenating the markdown texts, we generate collections with lengths of 64K, 128K, 512K, and 1M tokens.101010We can easily construct collections with arbitrary lengths exceeding 1M tokens.
@@
 From the 662 scientific articles across 16 clusters, we construct 2,612 article collections spanning all four length levels, including 2,027 generated through random sampling and 585 through graph traversal.
+3.2 Creating Databases and SQL Queries
+Once the article collections are assembled, we construct a database for each (see step II in Figure 1), focusing on a core subset of elements shared across all articles: titles, authors, and references.
+Based on these key elements, each database comprises three tables: articles, article-author, and citing-cited.
+The articles table contains metadata such as the title, reference count, and title word count.
+The article-author table captures author information, including names and their positions in the corresponding author list.
+The citing-cited table records citation relations among articles.
+A detailed description is provided in Appendix A.
+This information is obtained from Semantic Scholar or derived via simple preprocessing, such as splitting titles and counting words.
+Manual corrections were applied as needed to align with the full-text Markdown articles.
 | SQL Commands | |
@@
 | Aggregating | 20 | SELECT MAX(author_count) FROM articles |
-| Sorting | 27 | SELECT title_word_count FROM articles ORDER BY author_count ASC |
+| Sorting | 27 | SELECT title_wstatisticsord_count FROM articles ORDER BY author_count ASC |
 | Filtering | 107 | SELECT author_name FROM article_author WHERE author_position {author-position} |
@@
 | Relational Filtering | 20 | SELECT COUNT(*) FROM articles WHERE article_id NOT IN (SELECT article_id_citing FROM citing_cited) AND article_id IN (SELECT article_id_cited FROM citing_cited) |
-3.2 Creating Databases and SQL Queries
-Once the article collections are assembled, we construct a database for each (see step II in Figure 1).
-We focus on a core subset of elements shared across all scientific articles: titles, authors, and references.
-Based on these key elements, each database comprises three tables: articles, article-author, and citing-cited.
-The articles table contains metadata such as the title, reference count, and title word count.
-The article-author table captures author information, including names and their positions in the corresponding author list.
-The citing-cited table records citation relations among articles.
-A detailed description of these tables is provided in Appendix A.
-This information is obtained from Semantic Scholar or derived via simple preprocessing, such as splitting titles and counting words.
 The core SQL111111Structured Query Language (SQL) is a standardized language for managing and querying relational databases.
 commands summarized in Table 3 form the foundation for building SQL queries, often combined with operators (comparison, arithmetic, logical) to create more complex queries (since WHERE is used to filter data, it always works in conjunction with SQL operators).
-Aside from the basic SQL commands in Table 4 (Aggregating, Sorting, and Filtering), we define composite commands based on their combinations (i.e., Filtering+Aggregating and Filtering+Sorting).
+Aside from the SQL commands in Table 3 (Aggregating, Sorting, and Filtering), we define composite commands based on their combinations (i.e., Filtering+Aggregating and Filtering+Sorting).
 Using these, we manually create SQL query templates targeting different topics related to key elements of scientific articles, including Author Count, Author List, Reference Count, Title List, and Title Word Count.
@@
 We collectively refer to the SQL commands in Table 4 as information processing skills, since they test different information processing capabilities.
-We have various templates per skill, designed to be applicable across all article collections.
+We have various templates per skill, designed to be applicable across article collections.
 For each collection, we randomly select 10 templates and instantiate them with collection-specific values for all placeholders.
@@
 We use Qwen2.5-Coder-32B-Instruct (Hui et al., 2024) to convert SQL queries into natural language questions.
-To ensure queries and questions are meaning preserving, we validate each generated question by converting it back to SQL and verifying that both queries produce identical results when executed against the collection database.
+To ensure queries and questions are meaning-preserving, we validate each generated question by converting it back to SQL and verifying that both queries produce identical results when executed against the collection database.
 For each SQL-collection pair, we repeat this process up to 10 times; if no valid question is obtained, we discard the query for that collection.
-Using the prompts in Appendix B, we successfully generate natural language questions for 82.9% of SQL-collection pairs.
+Using the prompts in Appendix E, we successfully generate natural language questions for 82.9% of SQL-collection pairs.
 Following this, we obtain 2,121 test questions for evaluating models against the four context lengths defined in Section 3.1, and 19,543 instances for training (see Section 4.2).
@@
 For each context length, our test partition covers all information processing skills and question topics described in Section 3.2.
-The distribution of SQL commands and operators is provided in Appendix C.
+The distribution of SQL commands and operators is provided in Appendix A, Table 11.
 3.4 Data Quality Validation
@@
 | 1M | 614 | 46.3 | 18.6 | 63.6 |
+To validate SciTrek’s quality, we crowdsourced human annotations through Prolific.121212https://www.prolific.com Annotators (English native speakers from the US or the UK) were asked to answer natural language questions from a random sample of 120 instances, with 20 instances representing each of the six skills listed in Table 4.
+Crowdworkers answered questions using database tables rather than full-text articles, as documents spanning 1M tokens are impractical for humans to review.
+Each instance was annotated by three participants who were given the same instructions as those used for model testing with database tables as context (Figure 7, Appendix E).
+In each annotation session, crowdworkers were given three questions with different contexts.
+To ensure annotation quality, each session included a quality control question.
+On average, annotators spent about 5.5 minutes per session and were compensated above the UK living wage, at £12 per hour.
+We measured inter-annotator agreement and the alignment between database-derived answers (from SQL query execution) and human responses using exact match.
+Annotators were considered in agreement if two or more provided the same answer.
+Similarly, database answers were considered aligned with human responses if they matched the responses of two or more annotators.
+Table 6 reports average agreement and alignment across 120 instances, showing strong inter-annotator agreement and high consistency between database-derived and human answers.
+Manual inspection of all 120 questions given to annotators reveals that 3 of the 19 Relational Filtering questions were ambiguous, explaining the gap between alignment and agreement for this category in Table 6.
+While Qwen2.5-Coder-32B-Instruct generally converts SQL queries to natural language accurately, the complexity of Relational Filtering queries can introduce ambiguity.
 | Skill | Agree (%) | Align (%) |
@@
 | All | 88.3 | 83.3 |
-To validate the quality of SciTrek, we used Prolific121212https://www.prolific.com/ to crow-source annotations by asking human annotators to provide answers to our natural language questions.
-We randomly sampled 120 instances representing the six skills listed in Table 4 (20 instances per skill).
-Each instance was independently annotated by three annotators.
-Annotators were asked to answer the questions using database tables rather than full-text articles, as documents spanning 1M tokens are impractical for humans to review.
-The metadata in the database tables were sourced from Semantic Scholar or generated through simple preprocessing as explained in Section 3.1, with manual corrections applied when necessary to make them aligned with the collected full-text articles in Markdown.
-More annotation details are in Appendix D.
-We measured inter-annotator agreement and the alignment of answers obtained from executing SQL queries on the database with human responses, using exact match.
-Annotators were considered in agreement if two or more provided the same answer to a question.
-Similarly, our answer was considered aligned with the humans if it matched the answer of two or more annotators.
-Table 6 reports the average agreement and alignment across 120 instances, showing that human annotators largely agree with each other and that our database-based answers are highly consistent with human responses.
-| Full-text Articles | Database TablesΔ | ||||||||
-| Model | CLen | 64K | 128K | 512K | 1M | 64K | 128K | 512K | 1M |
+| Full-text Articles | Database Tables | ||||||||
+| Model | Context Size | F-64 | F-128 | F-512 | F-1024 | D-64 | D-128 | D-512 | D-1024 |
 | Qwen2.5-7B-Instruct-1M | 1M | 4.5 | 2.8 | 0.3 | 0.0 | 20.5 | 14.3 | 5.0 | 2.0 |
 | Qwen2.5-14B-Instruct-1M | 1M | 8.3 | 6.5 | 1.6 | 0.1 | 33.3 | 27.2 | 11.0 | 5.9 |
+| Qwen3-4B-Instruct-2507 | 256K | 2.1 | 7.2 | — | — | 25.9 | 16.9 | 6.9 | 2.8 |
+| Qwen3-4B-Thinking-2507 | 256K | 41.1 | 29.3 | — | — | 90.2 | 83.1 | 71.8 | 52.3 |
+| Qwen3-30B-A3B-Instruct-2507 | 256K | 5.4 | 3.2 | — | — | 29.8 | 21.2 | 6.0 | 4.2 |
+| Qwen3-30B-A3B-Thinking-2507 | 256K | 53.3 | 42.0 | — | — | 92.3 | 86.2 | 73.5 | 61.1 |
 | Gemma-3-27B-IT | 128K | 6.2 | 3.4 | — | — | 31.8 | 25.0 | 11.3 | 6.1 |
@@
 In our experiments, we evaluate both proprietary and open-weight models from various families that support contexts exceeding 128K and have shown strong performance on language understanding and mathematical reasoning benchmarks, e.g., MMLU (Hendrycks et al., 2021a) and MATH (Hendrycks et al., 2021b).
-For each family, we select the largest model we can feasibly run with our resources.131313All our experiments were conducted on 4 NVIDIA HGX H200s.
+For each family, we select the largest model we can feasibly run with our resources.131313Experiments were conducted on 4 NVIDIA HGX H200s.
 For example, we use Llama-4-Scout-17Bx16E-Instruct instead of Llama-4-Maverick-17B-128E-Instruct, and the distilled variant of DeepSeek-R1.
-Table 7 provides an overview of the models we consider, most of which were released within the past six months.
-For detailed model descriptions and settings, refer to Appendix F.
-As shown, the models vary in parameter scale and supported context length.
+Table 7 provides an overview of the models we consider, which vary in parameter scale and supported context length (most were released within the past six months).
+For detailed model descriptions and settings, refer to Appendix B.
 We evaluate models in two context settings: (1) using the full-text scientific articles within a given collection as context and (2) using only the corresponding database tables.
 All models in our study are instruction-tuned, and we employ a uniform set of prompts across them, as detailed in Appendix E.
-Models generate three answers for each question.
 We assess performance using average exact match and F1, as the expected outputs are factual items with minimal variation, such as specific numbers, author names, or article titles.
+For higher confidence, reported metrics are averages across three runs.
 4.1 Zero-shot Prompting
@@
 Perhaps unsurprisingly, proprietary models significantly outperform open-weight ones.
-We observe similar tendencies when using F1 as the evaluation metric (see Appendix G).
-Figure 2 presents a fine-grained analysis across question topics (e.g., author count, reference count, citation relations), skills (e.g., aggregating, sorting), and subjects (e.g., Economics, Biology) when using full-text articles as context.
-As can be seen, model performance shows little variation by subject.
-However, most models struggle with sorting tasks, while performing better on aggregation.
+We observe similar tendencies when using F1 as the evaluation metric (see Appendix C).
+Figure 2 presents a fine-grained analysis for the best six models across question topics (e.g., author count, reference count, citation relations), skills (e.g., aggregating, sorting), and subjects (e.g., Economics, Biology) when using full-text articles as context.
+Model performance shows little variation by subject.
+However, most models struggle more with sorting tasks, while performing better on aggregation.
 Performance is lowest on citation-related questions (i.e., Citation Relation, Reference Count), and somewhat higher on author- and title-related questions.
 4.2 Long-context Post-training
-Leveraging the data generation methodology described in Section 1, we curate a substantial training dataset (19,543 instances) across four context lengths (64K, 128K, 512K, and 1M).
+Leveraging the data generation methodology described in Section 1, we curate a substantial training dataset (19,543 instances) across four context lengths including 64K, 128K, 512K, and 1M.
 We use this data to assess whether more training could improve the performance of open-weight models on full-text articles.
@@
 For RL, we use GRPO (Shao et al., 2024) with a mixed reward of EM and F1 to encourage the model to produce both reasoning traces and answers.
-GRPO was chosen due to its success in similar verifiable and long-context tasks (Shao et al., 2024; Mroueh, 2025; Gurung & Lapata, 2025; Zheng et al., 2025).
-We found a simple
+GRPO was chosen due to its success in similar verifiable and long-context tasks (Shao et al., 2024; Mroueh, 2025; Gurung and Lapata, 2025; Zheng et al., 2025).
+We found a simple sum of EM and F1 to improve accuracy over time while still providing a useful training signal from difficult questions.
+The prompt for generating reasoning traces is provided in Appendix E.
+Since GRPO optimization is time-consuming, we restrict training to a single epoch which takes around 5 days.
 | Length | Topic | Skills | ||||
@@
 | o4-mini (ZS) | 48.4 | — | 62.5 | 22.3 | 53.7 | 39.3 |
-sum of EM and F1 to improve accuracy over time while still providing a useful training signal from difficult questions.
-The prompt for generating reasoning traces is provided in Appendix H.
-Since GRPO optimization is time-consuming, we restrict training to a single epoch which takes around 5 days.
 The results in terms of exact match in Table 8 demonstrate that both SFT and GRPO slightly improve performance across dimensions, surpassing GPT-4.1 despite its much larger parameter size.
-However, Qwen2.5-7B-Instruct-1M is still not able to generalize to longer inputs, but does improve on questions related to out-of-distribution question topics and information processing skills.
-5 Model Failure Analyses
-We conduct additional analyses to understand model failures on SciTrek when using full-text articles as context.
-Table 9 shows correlations between model performance and various factors: the input article count, the question length (LenQ), the length of the underlying SQL query (LenSQL), and the answer length (LenA).
-They indicate that model performance correlates primarily with the question length, while o4-mini shows correlations with the input article count and the SQL query length.
-| Model | #Articles | LenQ | LenSQL | LenA |
+However, Qwen2.5-7B-Instruct-1M is still not able to generalize to longer inputs, but does improve on out-of-distribution question topics and information processing skills.
+5 Model Failure Analysis
+| Model | #Articles | Question Length | SQL Length | Ground-truth Answer Length |
 | Qwen2.5-7B-1M | 0.04 | 0.15∗ | 0.11 | 0.01 |
 | Qwen2.5-14B-1M | 0.04 | 0.16∗ | 0.07 | 0.06 |
+| Qwen3-4B-Instruct-2507 | 0.01 | 0.16∗ | 0.05 | 0.03 |
+| Qwen3-4B-Thinking-2507 | 0.06 | 0.14∗ | 0.19∗ | 0.01 |
+| Qwen3-30B-A3B-Instruct-2507 | 0.02 | 0.15∗ | 0.06 | 0.08 |
+| Qwen3-30B-A3B-Think-2507 | 0.12∗ | 0.08 | 0.19∗ | 0.02 |
 | Gemma-3-27B-IT | 0.08 | 0.14∗ | 0.02 | 0.08 |
@@
 | Llama-3.3-70B | 0.06 | 0.16∗ | 0.03 | 0.05 |
-| DeepSeek-R1-Llama-70B | 0.15∗ | 0.16∗ | 0.02 | 0.08 |
+| DeepSeek-R1-Distill-Llama-70B | 0.15∗ | 0.16∗ | 0.02 | 0.08 |
 | Gemini 2.5 Pro | 0.16∗ | 0.22∗ | 0.03 | 0.01 |
@@
 | o4-mini | 0.14∗ | 0.09 | 0.25∗ | 0.06 |
-| Qwen2.5-7B-1M (SFT) | 0.03 | 0.32∗ | 0.09 | 0.14∗ |
-| Qwen2.5-7B-1M (GRPO) | 0.07 | 0.13∗ | 0.02 | 0.08 |
+| Qwen2.5-7B-Instruct-1M (SFT) | 0.03 | 0.32∗ | 0.09 | 0.14∗ |
+| Qwen2.5-7B-Instruct-1M (GRPO) | 0.07 | 0.13∗ | 0.02 | 0.08 |
+We conduct additional analysis to understand model failures on SciTrek when using full-text articles as context.
+Table 9 shows correlations between model performance and various factors: the number of input articles, the question length (LenQ), the length of the underlying SQL query (LenSQL), and the answer length (LenA).
+They indicate that model performance correlates primarily with the question length, while o4-mini shows correlations with the number of input articles and the length of the SQL query.
 Inspection of model answers reveals several failure patterns: (1) weaker models tend to respond by simply outputting “NULL”, which suggests they rely more on the instruction to produce a fallback answer rather than genuinely understanding the given context; (2) models often fail to follow the specified output formats, e.g., returning author lists when author numbers are requested; (3) models sometimes provide incomplete answers when lists are requested for questions that require aggregation; and (4) models misinterpret compound conditions, particularly struggling with negation-based filtering (e.g., the last example in Table 1).
-Detailed analyses are provided in Appendix I.1.
-We also manually analyzed the reasoning traces from our GRPO-based model (see examples in Appendix I.2).
-Specifically, we examined whether the reasoning is logically sound and individual steps obtain accurate information from specific contexts.
-We find that (1) the overall reasoning remains logically sound at 64K and 128K context lengths, although the model occasionally includes extraneous steps unrelated to the question; and (2) despite not being trained for 512K, the model still demonstrates coherent reasoning, but with reduced accuracy in specific steps compared to 64K and 128K; and (3) most reasoning steps that require counting are not accurate especially for references, which leads to incorrect answers.
-These findings indicate that GRPO-based reinforcement learning improves abstract reasoning, without enhancing the model’s accuracy in fine-grained operations, such as counting the number of references.
+More details of failure analysis are provided in Appendix D.
+We also manually analyzed 200 reasoning traces from Qwen2.5-7B-1M (GRPO), Qwen3-4B-A3B-Thinking-2507 and DeepSeek-R1-Distill-Llama-70B for 25 random SQL query templates.141414Reasoning traces from Qwen2.5-7B-1M (GRPO), Qwen3-4B-A3B-Thinking-2507, and DeepSeek-R1-Distill-Llama-70B are lengthy, averaging 638, 1,685, and 2,320 respectively.
+We will release these traces along with our code, but we omit them here due to space constraints.
+We cannot obtain reasoning traces from Gemini 2.5 Pro, GPT-4.1 and o4-mini for analysis.
+Specifically, we examined whether the reasoning is logically sound and whether individual steps obtain accurate information from specific contexts.
+We find that (1) the overall reasoning remains logically sound at 64K and 128K context lengths for all three models, although their traces occasionally include repeated steps or extraneous operations unrelated to the question; (2) despite not being trained for 512K, Qwen2.5-7B-1M (GRPO) still demonstrates coherent reasoning, but with reduced accuracy in specific steps compared to 64K and 128K; and (3) most reasoning steps that require counting are not accurate especially for references, which leads to incorrect answers from these models especially for Qwen2.5-7B-1M (GRPO).
+These findings indicate that GRPO-based reinforcement learning improves abstract reasoning without enhancing the model’s accuracy in fine-grained operations, such as counting references.
 6 Conclusion
-This paper introduced SciTrek, a benchmark designed for testing the ability of LLMs to perform multi-document information synthesis and structured reasoning over full-text scientific articles.
-By generating questions and ground-truth answers through a SQL backbone over article metadata, we provide a framework with explicit reasoning steps that is highly scalable and enables fine-grained error analysis.
-Our extensive evaluation demonstrates that SciTrek poses a significant challenge to both open-weight and proprietary LLMs, with only modest performance gains observed through supervised fine-tuning and reinforcement learning.
-Specifically, models struggle with compound logical conditions, fail at tasks requiring sorting, and often produce incomplete or badly formatted outputs.
-We believe the proposed methodology for creating SciTrek generalizes beyond titles, authors, and references to encompass broader metadata elements and domains, offering a tool to diagnose persistent shortcomings in LCLMs and improve their ability to support scientific workflows.
-Appendix A Description of SciTrek’s Database
+We introduce SciTrek, a benchmark for evaluating LLMs’ ability to perform multi-document information synthesis and structured reasoning over full-text scientific articles.
+By generating questions and ground-truth answers through a SQL backbone operating on article metadata, we provide a highly scalable framework with explicit reasoning processes that enables fine-grained error analysis.
+Our extensive evaluation demonstrates that SciTrek poses a significant challenge to open-weight and proprietary LLMs, with only modest performance gains observed through supervised fine-tuning and reinforcement learning.
+Specifically, models struggle with compound logical conditions, fail at sorting tasks, and often produce incomplete or poorly formatted outputs.
+We believe the proposed methodology for creating SciTrek generalizes beyond titles, authors, and references to encompass broader metadata elements and domains, offering a tool to diagnose persistent shortcomings in LCLMs and improve their capability to support scientific workflows.
+7 Limitations
+While SciTrek benchmarks fundamental reasoning capabilities (including filtering, sorting, aggregating and their combinations) of LCLMs on scientific articles, it covers only a limited set of core elements of scientific articles, specifically titles, authors, and references.
+Future work will explore additional long-context model capabilities for understanding scientific articles, such as interpreting figures and performing domain-specific content reasoning.
+Appendix A Details on SciTrek Construction
 We construct a database with three tables for each article collection, using metadata from Semantic Scholar along with basic preprocessing. A description of these tables is provided in Table 10.
-| Database Table | Column Name | Data Type | Description |
+Our test dataset covers all SQL commands and operators that are listed in Table 3. The detailed distribution is shown in Table 11.
+| Database | Column | Data | |
+| Table | Name | Type | Description |
 | articles | articleid | String | the unique identifier of the article |
@@
 | articleidcited | String | the identifier of the article that is cited by another article |
-Appendix B Prompts for Converting Between SQL Queries and Natural Language Questions
-We generate natural language questions by prompting Qwen2.5-Coder-32B to convert SQL queries to natural language.
-Our prompt is presented in Figure 3.
-The prompt that we use to convert natural language questions back to to SQL queries is shown in Figure 4.
-Appendix C Test Set Coverage of SQL Commands and Operators
-Our test dataset covers all SQL commands and operators that are listed in Table 4. The detailed distribution is shown in Table 11.
-| Command/Operator | #Instances | Proportion (%) |
-| SELECT | 2,121 | 100.00% |
-| WHERE | 1,821 | 85.86% |
-| 1,032 | 48.66% | |
-| IN | 682 | 32.15% |
-| OR | 616 | 29.04% |
-| ORDER BY | 591 | 27.86% |
-| 476 | 22.44% | |
-| 450 | 21.22% | |
-| COUNT | 370 | 17.44% |
-| ASC | 301 | 14.19% |
-| DESC | 287 | 14.19% |
-| DISTINCT | 280 | 13.20% |
-| 215 | 10.14% | |
-| MAX | 213 | 10.04% |
-| AND | 195 | 9.19% |
-| GROUP BY | 175 | 8.25% |
-| 175 | 8.25% | |
-| % | 164 | 7.73% |
-| 154 | 7.26% | |
-| NOT | 154 | 7.26% |
-| 153 | 7.21% | |
-| AVG | 135 | 6.36% |
-| MIN | 133 | 6.27% |
-| BETWEEN | 106 | 5.00% |
-| SUM | 102 | 4.81% |
-| 100 | 4.71% | |
-| LIKE | 50 | 2.36% |
-| + | 48 | 2.26% |
-| - | 23 | 1.08% |
-Appendix D Human Annotation Details
-We validated the quality of our curated questions and answers with a human annotation study.
-We recruited crowdworkers via Prolific (https://www.prolific.com/), specifically native English speakers from the US or UK.
-Our annotators were compensated above the UK living wage, at £12 per hour.
-We randomly sampled 120 instances from our test dataset (20 per skill in Table 4).
-Each instance was annotated by three participants who were given the same instructions as those used for model testing with database tables as context (Figure 6, Appendix E).
-In each annotation session, crowdworkers were given three questions with different contexts.
-To ensure annotation quality, each session included a quality control question.
-On average, annotators spent about 5.5 minutes per session for the three questions.
-Appendix E Prompts for Model Testing
-We employ two different prompts for (zero-shot) model evaluation.
-The prompt for assessing LCLM capabilities against full-text articles is presented in Figure 5, and the prompt for using the corresponding database tables as the context is shown in Figure 6.
-Appendix F Experimental Details
-The details of the models we use in our experiments are presented in Table 12. For all models, we use their default inference configuration for answer generation from their Huggingface repositories.
+| Command/ | Command/ | ||||
+| Operator | Count | % | Operator | Count | % |
+| SELECT | 2,121 | 100.00 | GROUP BY | 175 | 8.25 |
+| WHERE | 1,821 | 85.86 | 175 | 8.25 | |
+| 1,032 | 48.66 | % | 164 | 7.73 | |
+| IN | 682 | 32.15 | 154 | 7.26 | |
+| OR | 616 | 29.04 | NOT | 154 | 7.26 |
+| ORDER BY | 591 | 27.86 | 153 | 7.21 | |
+| 476 | 22.44 | AVG | 135 | 6.36 | |
+| 450 | 21.22 | MIN | 133 | 6.27 | |
+| COUNT | 370 | 17.44 | BETWEEN | 106 | 5.00 |
+| ASC | 301 | 14.19 | SUM | 102 | 4.81 |
+| DESC | 287 | 14.19 | 100 | 4.71 | |
+| DISTINCT | 280 | 13.20 | LIKE | 50 | 2.36 |
+| 215 | 10.14 | 48 | 2.26 | ||
+| MAX | 213 | 10.04 | 23 | 1.08 | |
+| AND | 195 | 9.19 |
+Appendix B Experimental Settings
+Configuration details for the models we use are presented in Table 12.
+For all models, we use their default inference configuration for answer generation from their Huggingface repositories.
 We include both reasoning models and non-reasoning models.
-DeepSeek-R1-Distill-Llama-70B, Gemini 2.5 Pro and o4-mini are reasoning models.
-In our experiments, o4-mini conduct reasoning with the default thinking effort medium, Gemini 2.5 Pro was given a thinking budget of 512 tokens, and DeepSeek-R1-Distill-Llama-70B was prompted with the reasoning prompt in Figure 7.
+Qwen3-4B-Thinking-2507, Qwen3-30B-A3B-Thinking-2507, DeepSeek-R1-Distill-Llama-70B, Gemini 2.5 Pro and o4-mini are reasoning models.
+In our experiments, o4-mini conducts reasoning with the default thinking effort medium, Gemini 2.5 Pro was given a thinking budget of 512 tokens, and Qwen3-4B-Thinking-2507, Qwen3-30B-A3B-Thinking-2507, and DeepSeek-R1-Distill-Llama-70B were prompted with the reasoning prompt in Figure 6.
 | Model | #Parameters | Open Weights | Context Length | Release Date |
@@
 | Qwen2.5-14B-Instruct-1M | 14B | ✓ | 1,010,000 | Jan 2025 |
+| Qwen3-4B-Instruct-2507 | 4B | ✓ | 262,144 | July 2025 |
+| Qwen3-4B-Thinking-2507 | 4B | ✓ | 262,144 | July 2025 |
+| Qwen3-30B-A3B-Instruct-2507 | 3Bx10E | ✓ | 262,144 | July 2025 |
+| Qwen3-30B-A3B-Thinking-2507 | 3Bx10E | ✓ | 262,144 | July 2025 |
 | Gemma-3-27B-IT | 27B | ✓ | 131,072 | Mar 2025 |
@@
 | o4-mini | — | ✗ | 200,000 | Apr 2025 |
-Appendix G Additional Results
-Table 13 presents additional results using F1 as the evaluation metric.
-Models are tested against full-text scientific articles and database tables as context.
-The input length averages only 1,980 tokens when database tables of are used as context across different lengths of full-text article collections.
-| Full-text Articles | Database TablesΔ | ||||||||
-| Model | CLen | 64K | 128K | 512K | 1M | 64K | 128K | 512K | 1M |
+| Full-text Articles | Database Tables | ||||||||
+| Model | Context Size | F-64 | F-128 | F-512 | F-1024 | D-64 | D-128 | D-512 | D-1024 |
 | Qwen2.5-7B-Instruct-1M | 1M | 8.7 | 7.0 | 1.1 | 0.0 | 39.2 | 39.3 | 28.7 | 21.2 |
 | Qwen2.5-14B-Instruct-1M | 1M | 20.6 | 17.5 | 7.4 | 1.0 | 58.2 | 53.9 | 44.8 | 34.8 |
+| Qwen3-4B-Instruct-2507 | 256K | 7.2 | 7.3 | — | — | 47.2 | 46.0 | 38.9 | 31.1 |
+| Qwen3-4B-Thinking-2507 | 256K | 50.2 | 42.3 | — | — | 92.7 | 85.8 | 79.4 | 63.6 |
+| Qwen3-30B-A3B-Instruct-2507 | 256K | 15.0 | 16.4 | — | — | 55.6 | 53.8 | 44.7 | 42.0 |
+| Qwen3-30B-A3B-Thinking-2507 | 256K | 62.5 | 55.1 | — | — | 94.4 | 88.3 | 80.8 | 74.5 |
 | Gemma-3-27B-IT | 128K | 22.6 | 14.7 | — | — | 52.8 | 52.5 | 47.8 | 40.7 |
@@
 | o4-mini | 195K | 74.0 | 63.8 | — | — | 97.2 | 89.2 | 87.0 | 86.5 |
-Appendix H Reasoning Traces for Reinforcement Learning
-The prompt used by GRPO to generate reasoning traces is given in Figure 7.
-Appendix I Details for Model Failure Analyses
-I.1 Analyses on Model Answers
-To gain deeper insights into model behavior, we analyzed the failure patterns exhibited in model-generated answers on SciTrek.
+Appendix C Additional Results Using F1
+Table 13 presents additional results using F1 as the evaluation metric.
+Models are tested against full-text scientific articles and database tables as context.
+The input length averages only 1,980 tokens when database tables of are used as context across different lengths of full-text article collections.
+Appendix D Details on Model Failure Analysis
+To gain deeper insight into model behaviour, we analyzed the failure patterns exhibited in model-generated answers on SciTrek.
 We randomly selected 60 instances in the input length of 128K that were hard for all zero-shot and supervised models (10 instances per skill in Table 4).
-We find that models tend to output “NULL” particularly on samples that require the skill of filtering.
-As shown in Table 14, most open-weight models tend to randomly output “NULL” for any sample.
-This indicates that these models rely more on the instruction to produce a fallback “NULL” answer rather than on genuine understanding of the provided context.141414Our testing instruction prompted models to generate “NULL” when they could not get the answer from the context.
+We find that models tend to output “NULL" particularly on samples that require the skill of filtering.
+As shown in Table 14, most open-weight models tend to randomly output “NULL" for any sample.
+This indicates that these models rely more on the instruction to produce a fallback “NULL” answer rather than on genuine understanding of the provided context.151515Our instruction prompted models to generate “NULL” when they could not get the answer from the context.
 In addition, we find that models fail to follow the specified format for the answers, especially for sorting-related questions, as shown in Table 15.
@@
 The results in Table 17 show that all models struggle with negation-based filtering.
-| Aggregation | Sorting | Filtering | Filtering | Filtering | Relational | |
-| Model | +Aggregation | +Sorting | Filtering | |||
+We further manually analyzed 200 reasoning traces from reasoning models including Qwen2.5-7B-1M (GRPO), Qwen3-4B-A3B-Thinking-2507 and DeepSeek-R1-Distill-Llama-70B for 25 random SQL query templates.
+We cannot obtain reasoning traces from Gemini 2.5 Pro, GPT-4.1 and o4-mini for analysis.
+For each SQL query template, we have different SQL queries and questions for different sizes of contexts, while the questions require the same reasoning capability.
+There are no reasoning traces when the input is 512K or 1M tokens for Qwen3-4B-Thinking-2507 and DeepSeek-R1-Distill-Llama-70B, as they do not support the input of 512K or 1M tokens.
+| Model | A | S | F | F+A | F+S | RF |
 | Qwen2.5-7B-Instruct-1M | 70 | 90 | 80 | 90 | 80 | 80 |
@@
 | Qwen2.5-7B-Instruct-1M (GRPO) | 0 | 0 | 0 | 0 | 0 | 0 |
-| Aggregation | Sorting | Filtering | Filtering | Filtering | Relational | |
-| Model | +Aggregation | +Sorting | Filtering | |||
+| Model | A | S | F | F+A | F+S | RF |
 | Qwen2.5-7B-Instruct-1M | 0 | 0 | 0 | 0 | 0 | 0 |
@@
 | Qwen2.5-7B-Instruct-1M (GRPO) | 0 | 0 | 0 | 10 | 0 | 0 |
-| Aggregation | Sorting | Filtering | Filtering | Filtering | Relational | |
-| Model | +Aggregation | +Sorting | Filtering | |||
+| Model | A | S | F | F+A | F+S | RF |
 | Qwen2.5-7B-Instruct-1M | 10 | 0 | 10 | 0 | 0 | 0 |
@@
 | Qwen2.5-7B-Instruct-1M (GRPO) | 20 | 0 | 30 | 10 | 20 | 0 |
-| Filtering | Filtering | Filtering | Relational | Average | |
-| Model | +Aggregation | +Sorting | Filtering | ||
+| Model | F+A | F+S | F | RF | Avg |
 | Qwen2.5-7B-Instruct-1M | 0 | 0 | 0 | 0 | 0 |
@@
 | Qwen2.5-7B-Instruct-1M (GRPO) | 0 | 33.3 | 8.3 | 3.2 | 8.3 |
-I.2 Example Reasoning Traces from GRPO-based Model
-We randomly sampled reasoning traces for three SQL query templates from the GRPO-based model.
-The reasoning traces for different sizes of contexts are presented in Table LABEL:tab:example_reasoning_traces.
-[…]
-[…]
-Question
-Context 64K |
-What is the smallest number of authors for articles that have more than 13 references? |
-[…]
-To solve this problem, we need to follow these steps:
-1. Identify articles with more than 13 references.
-2. Determine the number of authors in each of those articles.
-3. Find the smallest number of authors among these articles.
-Let’s go through each step in detail:
-### Step 1: Identify articles with more than 13 references
-From the provided articles, we need to count the references in each article:
-1. **”Best Constants in Young’s Inequality, Its Converse, and Its Generalization to More than Three Functions”**
-- References: 13 (listed in the reference section)
-2. **”Optical detection of strain and doping inhomogenieties in single layer MoS2”**
-- References: 62 (listed in the reference section)
-3. **”Ambipolar device simulation based on the drift-diffusion model in ion-gated transition metal dichalcogenide transistors”**
-- References: 65 (listed in the reference section)
-4. **”The Impact of AI in Physics Education: A Comprehensive Review from GCSE to University Levels”**
-- References: 32 (listed in the reference section)
-Only the articles titled ”Optical detection of strain and doping inhomogenieties in single layer MoS2”, ”Ambipolar device simulation based on the drift-diffusion model in ion-gated transition metal dichalcogenide transistors”, and ”The Impact of AI in Physics Education: A Comprehensive Review from GCSE to University Levels” have more than 13 references.
-### Step 2: Determine the number of authors in each of these articles
-Now, let’s count the authors in each of these articles:
-1. **”Optical detection of strain and doping inhomogenieties in single layer MoS2”**
-- Authors: A. Michail, N. Delikoukos, J. Parthenios, C. Galiotis, K. Papagelis (5 authors)
-2. **”Ambipolar device simulation based on the drift-diffusion model in ion-gated transition metal dichalcogenide transistors”**
-- Authors: Akiko Ueda, Yijin Zhang, Nobuyuki Sano, Hiroshi Imamura, Yoshihiro Iwasa (5 authors)
-3. **”The Impact of AI in Physics Education: A Comprehensive Review from GCSE to University Levels”**
-- Authors: Will Yeadon, Tom Hardy (2 authors)
-### Step 3: Find the smallest number of authors
-Among the articles with more than 13 references, the smallest number of authors is 2.
-[…]
-[…]
-[…]
-Question
-Context 128K |
-What is the smallest number of authors for articles that have more than 25 references? |
-[…]
-To solve this problem, we need to follow the steps outlined in the question. Let’s go through each step in detail.
-### Step 1: Identify Articles with More Than 25 References
-First, we need to go through each provided article and determine which ones have more than 25 references. We’ll count the words in the titles using spaces as delimiters to ensure accuracy.
-#### Article Titles and Reference Counts:
-1. **Article Title:** ”A SYSTEMATIC EVALUATION OF LARGE LANGUAGE MODELS OF CODE”
-- Words: Count spaces to determine words.
-A SYSTEMATIC EVALUATION OF LARGE LANGUAGE MODELS OF CODE
-Spaces: 8 words (A, SYSTEMATIC, EVALUATION, OF, LARGE, LANGUAGE, MODELS, OF, CODE)
-- Number of references: 27 (as given in the problem statement).
-This article has more than 25 references.
-2. **Article Title:** ”Reputation-based partner choice promotes cooperation in social networks”
-- Words: Count spaces to determine words.
-Reputation-based partner choice promotes cooperation in social networks
-Spaces: 9 words (Reputation-based, partner, choice, promotes, cooperation, in, social, networks).
-- Number of references: 26 (as given in the problem statement).
-This article has more than 25 references.
-<...we cut it down to save space...>
-### Step 2: Determine the Smallest Number of Authors in Articles with More Than 25 References
-Now that we’ve identified the articles with more than 25 references, let’s count the authors in each of these articles.
-1. **Article Title:** ”A SYSTEMATIC EVALUATION OF LARGE LANGUAGE MODELS OF CODE”
-- Authors: Frank F. Xu, Uri Alon, Graham Neubig, Vincent J. Hellendoorn (4 authors).
-2. **Article Title:** ”Reputation-based partner choice promotes cooperation in social networks”
-- Authors: Feng Fu, Christoph Hauert, Martin A. Nowak, Long Wang (4 authors).
-3. **Article Title:** ”Efficient Multimodal Sampling via Tempered Distribution Flow”
-- Authors: Yixuan Qiu, Xiao Wang (2 authors).
-4. **Article Title:** ”Ultralow contact resistance between semimetal and monolayer semiconductors”
-- Authors: Pin-Chun Shen, Cong Su, Yuxuan Lin, Ang-Sheng Chou, Chao-Ching Cheng, Ji-Hoon Park, Ming-Hui Chiu, Ang-Yu Lu, Hao-Ling Tang, Mohammad Mahdi Tavakoli, Gregory Pitner, Xiang Ji, Zhengyang Cai, Nannan Mao, Jiangtao Wang, Vincent Tung, Ju Li, Jeffrey Bokor, Alex Zettl, Chih-I Wu, Tomás Palacios, Lain-Jong Li, Jing Kong (11 authors).
-5. **Article Title:** ”Normal forms for singularities of pedal curves produced by non-singular dual curve germs in Sn”
-- Authors: Takashi Nishimura (1 author).
-### Step 3: Identify the Smallest Number of Authors
-[…]
-### Final Answer:
-[…]
-[…]
-[…]
-Question
-Context 512K |
-What is the smallest number of authors for articles that cite more than 84 references? |
-[…]
-To solve this problem, we need to follow these steps:
-1. Identify articles that cite more than 84 references.
-2. Determine the number of authors in each qualifying article.
-3. Identify the smallest number of authors among these articles.
-4. Provide the final answer in the required format.
-Let’s go through each step in detail.
-### Step 1: Identify articles that cite more than 84 references
-First, we will examine each provided article to determine which ones have more than 84 references.
-1. **Article Title:** ”Macroscopic model and statistical model to characterize electromagnetic information of a digital coding metasurface”
-Number of words in title: Counting spaces, we have ”Macroscopic model and statistical model to characterize electromagnetic information of a digital coding metasurface” → Words are: Macroscopic, model, and, statistical, model, to, characterize, electromagnetic, information, of, a, digital, coding, metasurface.
-(12 words).
-Number of references: 85 (as stated in the problem).
-2. **Article Title:** ”Worst-case VaR and robust portfolio optimization with interval random uncertainty set”
-Number of words in title: ”Worst-case VaR and robust portfolio optimization with interval random uncertainty set” → Words are: Worst-case, VaR, and, robust, portfolio, optimization, with, interval, random, uncertainty, set.
-(11 words).
-Number of references: 84 (as stated in the problem).
-<...we cut it down to save space...>
-Now we have identified the following articles that cite more than 84 references:
-- ”Macroscopic model and statistical model to characterize electromagnetic information of a digital coding metasurface” (85 references).
-- ”Worst-case VaR and robust portfolio optimization with interval random uncertainty set” (84 references).
-- ”Direct Transmission of Digital Message via Programmable Coding Metasurface” (82 references).
-- ”Bayesian experimental design for linear elasticity” (84 references).
-- ”Metaverse for Social Good: A University Campus Prototype” (84 references).
-- ”Alloying effect-induced electron polarization drives nitrate electroreduction to ammonia” (84 references).
-### Step 2: Determine the number of authors in each qualifying article
-Next, we will count the authors in each of these identified articles.
-1. **”Macroscopic model and statistical model to characterize electromagnetic information of a digital coding metasurface”**
-Authors listed: Rui Wen Shao, Wensheng Gan (listed as corresponding author), Philip S. Yu (listed as corresponding author).
-Number of authors: 3.
-2. **”Worst-case VaR and robust portfolio optimization with interval random uncertainty set”**
-Authors listed: Wei Chen, Shaohua Tan, Dongqing Yang.
-Number of authors: 3.
-3. **”Direct Transmission of Digital Message via Programmable Coding Metasurface”**
-Authors listed: Tie Jun Cui, Shuo Liu, Guo Dong Bai, Qian Ma.
-Number of authors: 4.
-4. **”Bayesian experimental design for linear elasticity”**
-Authors listed: Haihan Duan, Jiaye Li, Sizheng Fan, Zhonghao Lin, Xiao Wu, Wei Cai (listed as corresponding author).
-Number of authors: 6.
-<...we cut it down to save space...>
-### Step 3: Identify the smallest number of authors among these articles
-Among the identified articles that cite more than 84 references, the smallest number of authors is 3.
-### Final Answer
-[…]
-[…]
-[…]
-[…]
-Question
-Context 1M |
-What is the smallest number of authors for articles that cite more than 51 references? |
-[…]
-[…]
-[…]
-Question
-Context 64K |
-How many different articles has Shi Jin been listed as the first author of? |
-[…]
-To determine how many different articles Shi Jin has been listed as the first author of, we need to go through each article provided in the documents and check his position as an author.
-Let’s examine each article one by:
-1. **Broadband diffusion of terahertz waves by multi-bit coding metasurfaces**:
-- Authors: Li-Hua Gao, Qiang Cheng, Jing Yang, Shao-Jie Ma, Jie Zhao, Shuo Liu, Hai-Bing Chen, Qiong He, Wei-Xiang Jiang, Hui-Feng Ma, Qi-Ye Wen, Lan-Ju Liang, Biao-Bing Jin, Wei-Wei Liu, Lei Zhou, Jian-Quan Yao, Pei-Heng Wu, Tie-Jun Cui
-- Shi Jin is not listed as the first author here.
-2. **Large Intelligent Surface-Assisted Wireless Communication Exploiting Statistical CSI**:
-- Authors: Yu Han, Wankai Tang, Shi Jin, Chao-Kai Wen, Xiaoli Ma
-- Shi Jin is listed as the fourth author but not the first.
-3. **Intelligent Reflecting Surface Enhanced Wireless Network: Joint Active and Passive Beamforming Design**:
-- Authors: Qingqing Wu, Rui Zhang
-- Shi Jin is not listed as an author here.
-4. **Wireless Communication Based on Information Metasurfaces**:
-- Authors: Jun Yan Dai, Wankai Tang, Ming Zheng Chen, Chi Hou Chan, Qiang Cheng, Shi Jin, Tie Jun Cui
-- Shi Jin is listed as the seventh author but not the first.
-After examining all provided articles, Shi Jin has not been listed as the first author in any of them.
-Thus, the number of different articles Shi Jin has been listed as the first author of is .
-[…]
-[…]
-[…]
-Question
-Context 128K |
-How many distinct articles has Tang been the first author of? |
-[…]
-To solve this problem, we need to follow a systematic approach by analyzing each article provided in the documents. Let’s go through each step in detail.
-### Step 1: Identify Articles Written by Tang as the First Author
-First, we need to examine each article to determine if Tang is listed as the first author. We will go through each article one by:
-#### Article 1: ”The humping phenomenon during high speed gas metal arc welding”
-Authors listed: T. C. Nguyen, D. C. Weckman, D. A. Johnson, H. W. Kerr (D. C. Weckman is the corresponding author).
-Tang is not listed as an author in this article.
-#### Article 2: ”Physics of Humping Formation in Laser Powder Bed Fusion”
-Authors listed: C. Tang, K. Q. Le, C. H. Wong (C. Tang is the corresponding author).
-Tang is listed as the first author in this article.
-#### Article 3: ”Keyhole fluctuation and pore formation mechanisms during laser powder bed fusion additive manufacturing”
-Authors listed: Yuze Huang 1,2, Tristan G.
-Fleming 3, Samuel J.
-Clark 1,2,4, Sebastian Marussi 1,2, Kamel Fezzaa 4, Jeyan Thiyagalingam 5, Chu Lun Alex Leung 1,2, Peter D.
-Lee 1,2 (Yuze Huang is the first author).
-Tang is not listed as an author in this article.
-#### Article 4: ”Unveiling mechanisms and onset threshold of humping in high-speed laser welding”
-Authors listed: Zen-Hao Lai 1, Siguang Xu 2, Samuel J. Clark 3, Kamel Fezzaa 3, Jingjing Li 1,4 (Zen-Hao Lai is the first author).
-Tang is not listed as an author in this article.
-#### Article 5: ”Understanding of humping formation and suppression mechanisms using the numerical simulation”
-Authors listed: Dongsheng Wu a,b,c, Xueming Hua a,b,∗, Dingjian Ye a,b, Fang Li a,b (Dongsheng Wu is the first author).
-Tang is not listed as an author in this article.
-#### Article 6: ”Investigation of the humping formation in the high power and high speed laser welding”
-Authors listed: Yuewei Ai a,b, Ping Jia a,b,∗, Chunming Wang b, Gaoyang Mi b, Shaoning Geng a, Wei Liu a, Chu Han a (Yuewei Ai is the first author).
-Tang is not listed as an author in this article.
-### Step 2: Count Distinct Articles Written by Tang as the First Author
-From the analysis above, only one article lists Tang as the first author:
-- ”Physics of Humping Formation in Laser Powder Bed Fusion” by C. Tang, K. Q. Le, C. H. Wong.
-Thus, Tang has been the first author in exactly one distinct article.
-### Final Answer
-[…]
-[…]
-[…]
-Question
-Context 512K |
-How many different articles was Connor M. Brennan the first author of? |
-[…]
-To solve this problem, we need to follow the steps outlined in the question. Let’s go through each step carefully and determine the final answer.
-### Step 1: Identify Relevant Articles
-First, we need to scan through the provided articles to identify which ones have Connor M. Brennan as the first author. This will help us determine how many different articles he authored.
-#### Article Titles and Authors:
-1. **”The asymmetric effects of monetary policy on stock price bubbles”** by Christophe Blot, Paul Hubert, Fabien Labondance.
-2. **”Deconstructing monetary policy surprises: the role of information shocks”** by Marek Jarociński, Peter Karadi.
-3. **”Nominal Rigidities and the Dynamic Effects of a Shock to Monetary Policy”** by Lawrence J. Christiano, Martin Eichenbaum, Charles Evans.
-4. **”Risk Appetite and the Risk-Taking Channel of Monetary Policy”** by Michael D. Bauer, Ben S. Bernanke, Eric Milstein.
-5. **”The Macroeconomic Effects of the Federal Reserve’s Conventional and Unconventional Monetary Policies”** by Eric T. Swanson.
-6. **”Federal Reserve Information and the Behavior of Interest Rates”** by Christina D. Romer, David H. Romer.
-7. **”A Reassessment of Monetary Policy Surprises and High-Frequency Identification”** by Michael D. Bauer, Eric T. Swanson.
-8. **”An Alternative Explanation for the ’Fed Information Effect’”** by Michael D. Bauer, Eric T. Swanson.
-9. **”Monetary Policy Shocks: Data or Methods?”** by Connor M. Brennan, Margaret M. Jacobson, Christian Matthes, Todd B. Walker.
-10. **”High Asymmetric Effects of Monetary Policy on Stock Price Bubbles”** by Christophe Blot, Paul Hubert, Fabien Labondance (Note: This is a duplicate title and should be ignored).
-From this list, only one article has Connor M. Brennan as the first author:
-- **”Monetary Policy Shocks: Data or Methods?”** by Connor M. Brennan, Margaret M. Jacobson, Christian Matthes, Todd B. Walker.
-### Step 2: Verify Connor M. Brennan as the First Author
-Now, let’s confirm that Connor M. Brennan is indeed the first author of the identified article:
-- The article title is ”Monetary Policy Shocks: Data or Methods?”
-- The authors listed are Connor M. Brennan, Margaret M. Jacobson, Christian Matthes, Todd B. Walker.
-- Connor M. Brennan is listed first, confirming he is the first author.
-### Step 3: Count Words in the Title Using Spaces
-Next, we need to count the words in the title ”Monetary Policy Shocks: Data or Methods?” using spaces as delimiters. Let’s list each word:
-1. Monetary
-2. Policy
-3. Shocks:
-4. Data
-5. or
-6. Methods?
-There are 6 words in the title.
-### Final Answer
-[…]
-[…]
-[…]
-Question
-Context 1M |
-How many different articles has Saurabh Bodhe authored as the first author? |
-[…]
-[…]
-[…]
-Question
-Context 64K |
-How many different articles has Feng Miao authored? |
-[…]
-To determine how many different articles Feng Miao has authored, we need to go through each article provided and identify Feng Miao’s involvement in each one. Let’s analyze each article one by step.
-### Article 1: ”Hopping transport through defect-induced localized states in molybdenum disulphide”
-Authors listed: Hao Qiu, Tao Xu, Zilu Wang, Wei Ren, Haiyan Nan, Zhenhua Ni, Qian Chen, Shijun Yuan, Feng Miao, Fengqi Song, Gen Long, Yi Shi, Litao Sun, Jinlan Wang, Xinran Wang.
-Feng Miao is listed as an author in this article.
-### Article 2: ”Approaching the quantum limit in two-dimensional semiconductor contacts”
-Authors listed: Weisheng Li, Xiaoshu Gong, Zhihao Yu, Liang Ma, Wenjie Sun, Si Gao, Çağıl Köroğlu, Wenfeng Wang, Lei Liu, Taotao Li, Hongkai Ning, Dongxu Fan, Yifei Xu, Xuecou Tu, Tao Xu, Litao Sun, Wenhui Wang, Junpeng Lu, Zhenhua Ni, Jia Li, Xidong Duan, Peng Wang, Yuefeng Nie, Hao Qiu, Yi Shi, Eric Pop, Jinlan Wang, Xinran Wang.
-Feng Miao is not listed as an author in this article.
-### Article 3: ”Origins of genuine Ohmic van der Waals contact between indium and MoS2”
-Authors listed: Bum-Kyu Kim, Tae-Hyung Kim, Dong-Hwan Choi, Hanul Kim, Kenji Watanabe, Takashi Taniguchi, Heesuk Rho, Ju-Jin Kim, Yong-Hoon Kim, Myung-Ho Bae.
-Feng Miao is not listed as an author in this article.
-### Article 4: ”Van der Waals contacts between three-dimensional metals and two-dimensional semiconductors”
-Authors listed: Yan Wang, Jong Chan Kim, Ryan J. Wu, Jenny Martinez, Xiuju Song, Jieun Yang, Fang Zhao, Andre Mkhoyan, Hu Young Jeong, Manish Chhowalla.
-Feng Miao is not listed as an author in this article.
-### Conclusion
-Based on the analysis, Feng Miao has authored only one article, which is ”Hopping transport through defect-induced localized states in molybdenum disulphide.”
-[…]
-[…]
-[…]
-Question
-Context 128K |
-How many different articles has JIAYI SUN authored? |
-[…]
-To solve this problem, we need to follow these steps:
-1. Identify articles authored by JIAYI SUN.
-2. Count the number of distinct articles authored by JIAYI SUN.
-3. Provide the final answer in the required format.
-Let’s go through each step in detail:
-### Step 1: Identify articles authored by JIAYI SUN
-From the provided articles, we need to examine each one to determine if JIAYI SUN is listed as an author. Let’s review each article:
-1. **Article Title:** ”Solipsis: A Decentralized Architecture for Virtual Environments”
-Authors listed: Davide Frey, Jérôme Royan, Romain Piegay, Anne-Marie Kermarrec, Emmanuelle Anceaume, Fabrice Le Fessant, and Jiayi Sun.
-JIAYI SUN is present here.
-2. **Article Title:** ”Metaverse: Survey, Applications, Security, and Opportunities”
-Authors listed: Jiayi Sun, Wensheng Gan, Han-Chieh Chao, and Philip S. Yu.
-JIAYI SUN is present here.
-3. **Article Title:** ”Big Data Meets Metaverse: A Survey”
-Authors listed: Jiayi Sun, Wensheng Gan, Zefeng Chen, Junhui Li, and Philip S. Yu.
-JIAYI SUN is present here.
-4. **Article Title:** ”A Dynamic Resource Allocation Framework for Synchronizing Metaverse with IoT Service and Data”
-Authors listed: Yue Han, Dusit Niyato, Cyril Leung, Chunyan Miao, and Jiayi Sun.
-JIAYI SUN is present here.
-Thus, JIAYI SUN has authored four distinct articles.
-### Step 2: Count the number of distinct articles authored by JIAYI SUN
-From the analysis above, we have identified the following articles authored by JIAYI SUN:
-1. ”Solipsis: A Decentralized Architecture for Virtual Environments”
-2. ”Metaverse: Survey, Applications, Security, and Opportunities”
-3. ”Big Data Meets Metaverse: A Survey”
-4. ”A Dynamic Resource Allocation Framework for Synchronizing Metaverse with IoT Service and Data”
-[…]
-### Final Answer
-[…]
-[…]
-[…]
-Question
-Context 512K |
-How many different articles has Qiang Cheng authored? |
-[…]
-To solve this problem, we need to follow a systematic approach by analyzing each provided article to identify instances where Qiang Cheng is listed as an author. Let’s go through each step in detail.
-### Step 1: Identify Articles Mentions Qiang Cheng as an Author
-First, we will examine each article title and author list to determine which articles include Qiang Cheng as an author.
-1. **Article Title:** ”Intelligent Reflecting Surface Enhanced Wireless Network: Joint Active and Passive Beamforming Design”
-[…]
-2. **Article Title:** ”Broadband continuous beam-steering with time-modulated metasurfaces in the near-infrared spectral regime”
-[…]
-3. **Article Title:** ”Path Loss Modeling and Measurements for Reconfigurable Intelligent Surfaces in the Millimeter-Wave Frequency Band”
-[…]
-4. **Article Title:** ”Index Modulation for 6G Wireless Networks Using Time-Varying Metasurfaces”
-[…]
-From this analysis, we see that Qiang Cheng is an author in four different articles.
-### Step 2: Count the Number of Articles Featuring Qiang Cheng
-Now that we have identified the articles where Qiang Cheng is an author, we can count them. The list of articles includes:
-1. ”Intelligent Reflecting Surface Enhanced Wireless Network: Joint Active and Passive Beamforming Design”
-2. ”Broadband continuous beam-steering with time-modulated metasurfaces in the near-infrared spectral regime”
-3. ”Path Loss Modeling and Measurements for Reconfigurable Intelligent Surfaces in the Millimeter-Wave Frequency Band”
-4. ”Index Modulation for 6G Wireless Networks Using Time-Varying Metasurfaces”
-[…]
-### Final Answer
-[…]
-[…]
-[…]
-Question
-Context 1M |
-How many different articles has A. N. Pettitt contributed to? |
-[…]
+Appendix E Prompts used for SciTrek Creation and Model Evaluation
+We generate natural language questions by prompting Qwen2.5-Coder-32B to convert SQL queries to natural language.
+Our prompt is presented in Figure 3.
+The prompt that we use to convert natural language questions back to to SQL queries is shown in Figure 4.
+We employ two different prompts for (zero-shot) model evaluation.
+The prompt for assessing LCLM capabilities against full-text articles is presented in Figure 5, and the prompt for using the corresponding database tables as context is shown in Figure 7.
+The prompt used by GRPO to generate reasoning races is given in Figure 6.